### v1

# Learning Operators through Coefficient Mappings in Fixed Basis Spaces

Abstract
Operator learning has emerged as a powerful paradigm for approximating solution operators of partial differential equations (PDEs) and other functional mappings. Classical approaches typically adopt a pointwise-to-pointwise framework, where input functions are sampled at prescribed locations and mapped directly to solution values. We propose the Fixed-Basis Coefficient to Coefficient Operator Network (FB-C2CNet), which learns operators in the coefficient space induced by prescribed basis functions. In this framework, the input function is projected onto a fixed set of basis functions (e.g., random features or finite element bases), and the neural operator predicts the coefficients of the solution function in the same or another basis. By decoupling basis selection from network training, FB-C2CNet reduces training complexity, enables systematic analysis of how basis choice affects approximation accuracy, and clarifies what properties of coefficient spaces (such as effective rank and coefficient variations) are critical for generalization. Numerical experiments on Darcy flow, Poisson equations in regular, complex, and high-dimensional domains, and elasticity problems demonstrate that FB-C2CNet achieves high accuracy and computational efficiency, showing its strong potential for practical operator learning tasks.
keywords:
Operator Learning , Random Feature Method, Function Encoders , Basis Functions
journal: Computer Methods in Applied Mechanics and Engineering
[…]
11footnotetext: Author names are ordered alphabetically to indicate equal contribution.
1 Introduction
Deep learning has recently emerged as a powerful paradigm in scientific computing, with applications ranging from solving differential equations to accelerating large-scale simulations and data-driven modeling [1, 2, 3, 4]. Among various methodologies, operator learning has garnered significant attention, as it provides a principled framework for approximating mappings between infinite-dimensional function spaces directly. Unlike neural PDE solvers such as PINNs [5], DeepRitz [6], and WANs [7], which approximate the solution of a partial differential equation (PDE) at finitely many spatial or temporal points, operator learning  [8, 9, 10] aims to approximate the entire mapping from input functions—such as initial conditions, boundary conditions, or source terms—to output functions, typically the corresponding PDE solutions. For instance, consider the operator $G:(a,f)\mapsto u,$ which maps a coefficient function $a$ and a source term $f$ to the solution $u$ of the elliptic PDE $\nabla\cdot(a\nabla u)=f,$ subject to appropriate boundary conditions. This example will serve as a recurring case study throughout this paper. This operator-based perspective represents a paradigm shift: once trained, operator learning models can serve as fast and generalizable solvers applicable to a wide range of problem settings. This stands in contrast to conventional numerical algorithms, which typically rely on case-specific discretizations and require repeated, computationally intensive simulations for each new instance.
Despite its great potential, operator learning faces several fundamental challenges. One of the most critical issues is the discretization of infinite-dimensional function spaces. In other words, a key question is: what should serve as the input and output representations of the neural network? To make the problem computationally tractable, these infinite-dimensional spaces must be approximated by finite-dimensional subspaces. In most existing operator learning frameworks [8, 9, 10, 11, 12, 13], functions are discretized by their nodal values at a collection of collocation points, which are then used as the neural network input and output. While this pointwise encoding is straightforward and widely adopted, it often suffers from drawbacks such as mesh dependence and deteriorating generalization as the resolution increases.
In this work, we propose a fixed basis coefficient-to-coefficient (FB-C2C) operator learning framework. The key idea is to represent both input and output functions in terms of carefully chosen basis functions, and to train a neural network to learn the mapping between their expansion coefficients. Let $G:\mathcal{X}\rightarrow\mathcal{Y}$ be an operator that maps an input function $f\in\mathcal{X}$ to an output function $G(f)\in\mathcal{Y}$. We approximate the input function space $\mathcal{X}$ by a finite-dimensional subspace $V_{\mathrm{in}}=\text{span}\{\phi_{j}(x)\}_{j=1}^{m_{1}}$ and the output function space $\mathcal{Y}$ by $V_{\mathrm{out}}=\text{span}\{\psi_{j}(y)\}_{j=1}^{m_{2}}$. For each input function $f(x)$, we approximate it as
$f(x)\approx\sum_{j=1}^{m_{1}}a_{j}\phi_{j}(x),$
and use the coefficient vector $\bm{a}=(a_{1},a_{2},\dots,a_{m_{1}})$ as the input to a neural network $NN_{\theta}$, where $\theta$ denotes the trainable parameters. The network outputs the coefficient vector corresponding to the output basis functions, yielding an approximation of the operator output as
$G(f)(y)\approx\sum_{j=1}^{m_{2}}\big[NN_{\theta}(\bm{a})\big]_{j}\,\psi_{j}(y).$
In this way, the proposed framework learns the mapping directly between input and output coefficients, thereby realizing a fixed basis coefficient-to-coefficient (FB-C2C) operator learning paradigm that provides a compact representation and facilitates improved accuracy and generalization compared to nodal-value encodings. The main idea of our method is illustrated in Figure 1. A detailed description of the proposed methodology will be presented in Section 2.
Figure 1: Fixed-Basis Coeffiecient to Coeffiecient Operator Network (FB-C2CNet) framework.
This section proceeds with a review of the relevant literature, followed by a summary of the main contributions of this work.
1.1 Related work
In recent years, leveraging neural network architectures for solving partial differential equations (PDEs) has attracted substantial attention in the scientific computing community. Such approaches are particularly promising in handling problems defined on complex domains and in incorporating empirical data into mathematical models. From a theoretical standpoint, neural networks have the capacity to overcome the curse of dimensionality in high-dimensional PDE problems [14, 15, 16, 17, 18]. Broadly speaking, existing neural network–based methods for solving PDEs can be categorized into two major classes.
The first is PDE solution approximation, where neural networks are trained to approximate the solution of a PDE directly. Representative approaches in this category include Physics-Informed Neural Networks (PINNs) [5, 19, 20], the Deep Ritz Method [6], Weak Adversarial Networks (WANs)  [7, 21], and random feature models [22, 23, 24, 13]. These methods typically design specialized loss functions that encode the underlying PDE constraints, boundary conditions, or variational formulations, thereby guiding the neural network toward solutions consistent with the governing equations. Such approaches have demonstrated notable success in forward problems, inverse problems, and data assimilation tasks. However, their efficiency often deteriorates as the dimensionality of the problem grows, and training may become prohibitively expensive or unstable for highly nonlinear systems.
The second class is operator learning, which shifts the perspective from approximating a single PDE solution to learning the underlying solution operator. In this framework, the goal is to approximate mappings from input functions—such as coefficients, source terms, or boundary conditions—to corresponding solution functions. This operator-centric viewpoint not only enables rapid solution of new problem instances once the operator is learned, but also facilitates generalization across families of PDEs. Representative methods include DeepONet [8] and the Fourier Neural Operator (FNO) [9], which introduced new architectures for handling function-to-function mappings efficiently. Building upon these seminal works, a growing number of operator-learning–based approaches have been developed to approximate increasingly complex classes of operators, ranging from multi-scale operators [25], to domain decomposition operators [26], to operators arising in physical and material sciences [27, 28]. These developments underscore the growing importance of operator learning as a paradigm that bridges classical numerical analysis with modern machine learning. Our work is centered on operator learning, and in what follows, we review existing literature and approaches that are closely connected to the present study.
Operator Learning
DeepONet [8] is one of the earliest and most influential works on operator learning. Building on the Universal Approximation Theorem for Operators, it introduces the framework:
$G(f)(y):=NN_{brach}(f)\cdot NN_{trunk}(y)=\underbrace{\sum_{k=1}^{p}\sum_{i=1}^{q}c_{i}^{k}\,\sigma\!\left(\sum_{j=1}^{m}\xi_{ij}^{k}\,f(x_{j})+\theta_{i}^{k}\right)}_{\text{Branch}}\;\underbrace{\sigma\!\left(\mathbf{w}_{k}\cdot\mathbf{y}+\zeta_{k}\right),}_{\text{Trunk}}$ (1)
where the Branch Net takes as input the function values at sensor points $f(x_{j})$, while the Trunk Net takes spatial coordinates and provides basis functions. The final output is obtained by combining the coefficients from the Branch Net with the basis from the Trunk Net. Following this, POD-DeepONet [10] replaced the Trunk Net with Proper Orthogonal Decomposition (POD) modes extracted from training data, while keeping the Branch Net to learn their coefficients. Later studies further explored alternative trunk representations, including fixed bases from Extreme Learning Machines (ELM) [11], finite element shape functions in Mesh-Informed Neural Networks (MINNs) [12], and random feature models with partition-of-unity strategies [13]. Despite these variations, the Branch Net still relies on $f(x_{j})$ as input. However, this design suffers from scalability issues: in higher dimensions, the number of sensor points becomes large, leading to high-dimensional inputs, larger networks, slower training, and potentially weaker generalization.
To address the challenges posed by high-dimensional inputs in operator learning, recent advances have explored integrating neural networks with principled dimensionality reduction strategies [29, 30, 31, 32, 33, 34]. Among these, three works are particularly relevant to our approach. The Basis Operator Network [35] represents input functions using basis coefficients: the Branch Net takes as input the coefficients of the selected basis functions, while the Trunk Net encodes spatial locations, with orthogonality of the Trunk Net preserved throughout training. A related idea is pursued in Basis-to-Basis Operator Learning Using Function Encoders [36], where the Trunk and Branch Nets are trained separately. In this framework, the trained Trunk Net is treated as a function encoder, and its output coefficients are subsequently used as inputs to the Branch Net. In the Resolution-Independent Neural Operator paper [37], the trunk network selects basis functions from a predefined dictionary of functions, and two dictionary learning algorithms are employed to adaptively learn a set of suitable continuous basis functions. The main limitation of these approaches lies in the fact that their basis functions must be obtained through training, which is both time-consuming and prone to instability. Our proposed approach is to directly select a function space—such as finite element basis functions or random feature models—as the function encoder, and then compute the coefficients of the input function with respect to these basis functions using the least squares method. This makes our method more effective, leading to shorter overall training time and improved stability. The coefficients are subsequently used as the input to the neural operator.
1.2 Contributions.
In this work, we propose the Fixed-Basis Coefficient Operator Network (FB-C2CNet), a new operator learning framework built upon fixed basis function systems. Our main contributions are summarized as follows:
  • 1.
    We introduce FB-C2CNet, which decouples operator learning into two stages: projection onto a fixed basis and learning in the coefficient space. This design eliminates the need to train the basis or encoder, thereby significantly reducing training cost.
  • 2.
    We investigate different choices of approximation function spaces, including random feature models and finite element bases, and further discuss strategies for computing the coefficients in a manner that facilitates learning by the neural operator. In particular, we analyze which characteristics of the coefficient representations are most critical for generalization, highlighting the role of effective rank and coefficient variations in producing representations that are both expressive and stable.
  • 3.
    We demonstrate the effectiveness of the proposed approach through extensive numerical experiments, including Darcy flow in 1D/2D regular domains, Poisson equations in complex domains, the Elastic equation, Darcy flow in complex domains, and high-dimensional Poisson equations, showing that our method achieves both high efficiency and accuracy.
  • 4.
    We further generalize the framework to multi-input/output operator learning, involving mappings between multiple input and output functions. We compare scalar-valued representations, obtained by concatenating input/output coefficients, with vector-valued constructions that employ shared vector-valued basis functions defined over all components. Numerical experiments show that the RFM-based vector-valued formulation consistently achieves higher stability and accuracy.
The remainder of the paper is organized as follows. Section 2 provides a detailed introduction to the proposed fixed basis coefficient-to-coefficient operator learning framework. Section 3 discusses the design choices of key components, including the selection of basis function spaces and strategies for coefficient computation. Section 4 presents numerical experiments, covering Darcy flow, Poisson equations on both regular and complex domains, as well as high-dimensional settings, and elastic plate equation to demonstrate the effectiveness, efficiency, and accuracy of our method. Section 5 concludes with a discussion and outlines directions for future work.
2 Methodology
In this work, our goal is to approximate the mapping $G:\mathcal{X}\mapsto\mathcal{Y}$ using neural networks, where $\mathcal{X}$ and $\mathcal{Y}$ denote infinite-dimensional spaces of real-valued functions defined on a bounded domain in $\mathbb{R}^{d}$. The operator $G$ maps an input function $f\in\mathcal{X}$ to the corresponding solution $u\in\mathcal{Y}$ of a given partial differential equation (PDE). Our approach focuses on the representation of the input function $f$. Instead of adopting a pointwise discretization, we employ a basis-function representation: the input function $f$ is encoded by its expansion coefficients with respect to a chosen set of basis functions, and the neural network is trained to predict the coefficients associated with the basis representation of the solution. Compared with pointwise discretization, using coefficients as the input encoding significantly reduces the dimensionality of the neural network input. This reduction in dimension decreases the number of network parameters and leads to faster convergence during training, particularly in two- or higher-dimensional settings. In addition, the coefficient-based representation provides a more compact and structured description of the function space, mitigates the redundancy inherent in dense pointwise sampling, improves numerical stability, and enhances interpretability by directly linking the learned coefficients to the underlying basis functions. It is worth emphasizing that the basis functions in our framework are pre-selected and fixed, rather than learned during training. This design makes the method more efficient and results in faster and more stable training.
The remainder of this section provides a detailed introduction to our method and elaborates on its specific components.
2.1 Problem formulation.
Let $\mathcal{X},\mathcal{Y}$ be function spaces, our goal is to use neural network to learn the operator
$G:\mathcal{X}\to\mathcal{Y}.$ (2)
Suppose we are given the training data in the form of input-output pair $\{f^{(k)},u^{(k)}\}_{k=1}^{M}\subset\mathcal{X}\times\mathcal{Y}$, where $f^{(k)}\in\mathcal{X}$ is usually the source term or initial condition for the PDE problem, and $u^{(k)}\in\mathcal{Y}$ is the solution of the PDE. To numerically resolve the PDE, a set of collocation points must be specified within the domain. The input functions are evaluated at the points $(\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{n_{1}})$ and the output functions at $(\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{y}_{n_{2}})$, respectively. Thus, the training data set becomes: $\{\mathbf{f}^{(k)}=[f^{(k)}(\mathbf{x}_{1}),f^{(k)}(\mathbf{x}_{2}),\dots,f^{(k)}(\mathbf{x}_{n_{1}})]^{T},\mathbf{u}^{(k)}=[u^{(k)}(\mathbf{y}_{1}),u^{(k)}(\mathbf{y}_{2}),\dots,u^{(k)}(\mathbf{y}_{n_{2}})]^{T}\}_{k=1}^{M}$. In practice, we approximate this mapping using a neural network $NN_{\bm{\theta}}$, where $\bm{\theta}$ represents the neural network’s parameters. In the following, we introduce our proposed method, termed FB-C2CNet.
2.2 Fixed-Basis Coefficient to Coefficient Operator Network (FB-C2CNet)
Our method consists of two main stages. The first stage is fixed basis function selection. In this stage, we specify two basis function systems, denoted as $V_{\text{in}}=\{\phi_{i}(\mathbf{x})\}_{i=1}^{m_{1}}$ and $V_{\text{out}}=\{\psi_{i}(\mathbf{y})\}_{i=1}^{m_{2}}$, which are employed to approximate the input function space $\mathcal{X}$ and the output function space $\mathcal{Y}$, respectively. In the subsequent step, a given input function $f(\mathbf{x})\in\mathcal{X}$ is represented in terms of the basis $V_{\text{in}}=\{\phi_{i}(\mathbf{x}),\phi_{2}(\mathbf{x}),\dots,\phi_{m_{1}}(\mathbf{x})\}$, i.e.,
$f(\mathbf{x})\approx\sum_{j=1}^{m_{1}}a_{j}\,\phi_{j}(\mathbf{x}),\quad\phi_{j}\in V_{\text{in}},f(\mathbf{x})\in\mathcal{X},$ (3)
where $\mathbf{a}=[a_{1},a_{2},\dots,a_{m_{1}}]\in\mathbb{R}^{m_{1}}$ are the expansion coefficients associated with $f$. The expansion coefficients $\mathbf{a}$ are obtained by solving the following regularized least-squares problem:
$\min_{\mathbf{a}\in\mathbb{R}^{m_{1}}}\;\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\left|f(\mathbf{x}_{i})-\sum_{j=1}^{m_{1}}a_{j}\,\phi_{j}(\mathbf{x}_{i})\right|^{2}+\lambda\|\mathbf{a}\|_{2}^{2},$ (4)
where $\{x_{i}\}_{i=1}^{n_{1}}$ are the sampling points in the input domain, and the second term provides $L^{2}$-regularization to stabilize the coefficient estimation. In the later sections of this paper, we will provide a more detailed discussion on the computation of the coefficients and analyze how different solution strategies may influence the performance of FB-C2CNet.
Following the above procedure, we obtain the coefficient representations $\{\mathbf{a}^{(k)}=[a^{(k)}_{1},a^{(k)}_{2},\dots,a^{(k)}_{m_{1}}]\}_{k=1}^{M}$, where each $\mathbf{a}^{(k)}\in\mathbb{R}^{m_{1}}$ corresponds to the expansion coefficients of the input function $f^{(k)}$ with respect to the input basis function systems $V_{\text{in}}$. We then take these coefficient representations as the input to our neural operator. This constitutes the second stage of our framework, in which a neural network is employed to approximate the underlying operator that maps the input coefficients $\mathbf{a}^{(k)}\in\mathbb{R}^{m_{1}}$ to the corresponding output coefficients $\mathbf{b}^{(k)}\in\mathbb{R}^{m_{2}}$ associated with the output basis function systems $V_{\text{out}}$. Specifically, the neural network $NN_{\bm{\theta}}$ is trained to approximate the mapping between the input and output coefficient representations, i.e.,
$NN_{\bm{\theta}}:\mathbf{a}\mapsto\mathbf{b},$ (5)
where $\mathbf{a}\in\mathbb{R}^{m_{1}}$ and $\mathbf{b}\in\mathbb{R}^{m_{2}}$ denote the coefficient representations of the input and output functions with respect to the fixed bases $V_{\text{in}}$ and $V_{\text{out}}$, respectively. In practice, we adopt a fully connected neural network (FNN) as the architecture for the operator approximation. Therefore, the overall structure of our proposed FB-C2CNet can be summarized as follows:
[…]
Here the set of trainable parameters in FB-C2CNet is denoted by $\bm{\theta}=\{c_{i}^{k},\,\xi_{ij}^{k},\,\eta_{i}^{k}\}$, which collectively define the weights and biases of the underlying fully connected network. $m_{1}$ denotes the number of basis functions in $V_{\text{in}}$, while $m_{2}$ denotes the number of basis functions in $V_{\text{out}}$. In the following, we present the training procedure adopted for FB-C2CNet.
Training method
The loss function employed to train FB-C2CNet is the relative loss, defined as
$L(\theta)=\frac{1}{M}\sum_{k=1}^{M}\frac{\left\|NN_{\theta}(\mathbf{a}^{(k)})\cdot\mathbf{\Psi}-\mathbf{u}^{(k)}\right\|_{2}}{\left\|\mathbf{u}^{(k)}\right\|_{2}},$ (7)
where $\{\mathbf{a}^{(k)}=[a_{1}^{(k)},a_{2}^{(k)},\dots,a_{m_{1}}^{(k)}]\}_{k=1}^{M}$ denotes the set of input coefficient vectors, $\{\mathbf{u}^{(k)}=[u^{(k)}(\mathbf{y}_{1}),u^{(k)}(\mathbf{y}_{2}),\dots,u^{(k)}(\mathbf{y}_{n_{2}})]\}_{k=1}^{M}$ represents the corresponding sampled outputs, and $\mathbf{\Psi}\in\mathbb{R}^{n_{2}\times m_{2}}$ is the basis evaluation matrix with entries
$\Psi_{ij}=\psi_{j}(\mathbf{y}_{i}),\quad 1\leq i\leq n_{2},\;1\leq j\leq m_{2}.$ (8)
Here, $NN_{\bm{\theta}}$ is the neural operator parameterized by $\bm{\theta}$.
We adopt the mean relative error loss instead of the plain mean squared error (MSE) in order to normalize the approximation error with respect to the magnitude of the ground truth. This normalization ensures scale-invariance across different training samples: functions with larger amplitudes do not dominate the optimization process, while functions with smaller amplitudes are not overlooked. Such a formulation is particularly important in operator learning, where the outputs $\{\mathbf{u}^{(k)}\}_{k=1}^{M}$ may vary significantly in scale across different samples. By employing the relative loss, FB-C2CNet achieves a more balanced training process and yields a more robust generalization performance. The overall procedure is illustrated in Figure 2 and Algorithm.
Figure 2: Workflow of the proposed Fixed-Basis Coefficient to Coefficient Network (FB-C2CNet), exemplified via RFM Basis functions. It consists of three main stages: (i) fixed basis function selection, (ii) coefficient representation of input and output functions, and (iii) coefficient-to-coefficient operator learning via a neural network.
Algorithm Fixed Space Coefficient-to-Coefficient Operator Learning
Phase I: Choose fixed input/output basis systems: $V_{\mathrm{in}}=\text{span}\{\phi_{j}(\bm{x})\}_{j=1}^{m_{1}}$,  $V_{\mathrm{out}}=\text{span}\{\psi_{j}(\bm{y})\}_{j=1}^{m_{2}}$.
Strategy 1: Random feature model
Input: $m_{1},m_{2}$, $R_{m}^{\mathrm{in/out}}$, $M_{p}^{\mathrm{in/out}},J_{n}^{\mathrm{in/out}}$
Procedure: 1. For $n=1,\dots,M_{p}^{\mathrm{in}}$ and $j=1,\dots,J_{n}^{\mathrm{in}}$: $\phi_{nj}(\bm{x})=\omega_{n}(\bm{x})\sigma(\bm{k}^{\mathrm{in}}_{nj}\!\cdot\!\bm{x}+b^{\mathrm{in}}_{nj}),$ with $\bm{k}^{\mathrm{in}}_{nj}\!\sim\!U([-R_{m}^{\mathrm{in}},R_{m}^{\mathrm{in}}]^{d})$, $b^{\mathrm{in}}_{nj}\!\sim\!U([-R_{m}^{\mathrm{in}},R_{m}^{\mathrm{in}}])$. 2. Similarly construct $\psi_{nj}(\bm{y})$ for output basis with $\bm{k}^{\mathrm{out}}_{nj}$, $b^{\mathrm{out}}_{nj}$. Output: $V_{\mathrm{in}}=\text{span}\{\phi_{j}(\bm{x})\}_{j=1}^{m_{1}}$, $V_{\mathrm{out}}=\text{span}\{\psi_{j}(\bm{y})\}_{j=1}^{m_{2}}$.
Strategy 2: Finite element model
Input: element types $E^{\mathrm{in/out}}$, polynomial degrees $p^{\mathrm{in/out}}$, evaluation points $\bm{\xi}^{\mathrm{in/out}}$
Procedure: 1. Generate the mesh $\mathcal{T}_{h}$; then, for the chosen element family and degree, construct the global FE basis. Output: $V_{\mathrm{in}}=\text{span}\{\phi_{j}(\bm{x})\}_{j=1}^{m_{1}}$, $V_{\mathrm{out}}=\text{span}\{\psi_{j}(\bm{y})\}_{j=1}^{m_{2}}$.
Phase II: Encode input coefficients via least squares.
Input:data sets $\{\mathbf{f}^{(k)}\}_{k=1}^{M}$, each $\mathbf{f}^{(k)}=\{f^{(k)}(\bm{x}_{i})\}_{i=1}^{n_{1}}$; penalty $\lambda$
for $k=1$ to $M$ do
  compute $\bm{a}^{(k)}\in\mathbb{R}^{m_{1}}$ by
   $\bm{a}^{(k)}=\underset{\bm{a}\in\mathbb{R}^{m_{1}}}{\arg\min}\;\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\left|f^{(k)}(\bm{x}_{i})-\sum_{j=1}^{m_{1}}a_{j}\phi_{j}(\bm{x}_{i})\right|^{2}+\lambda\|\bm{a}\|_{2}^{2}.$
end for
[…]
Phase III: Train coefficient-to-coefficient network $NN_{\theta}$.
Input:targets $\{\mathbf{u}^{(k)}\}_{k=1}^{M}$, each $\mathbf{u}^{(k)}=\{u^{(k)}(\bm{y}_{i})\}_{i=1}^{n_{2}}$; step size $\eta$
form $\Psi\in\mathbb{R}^{n_{2}\times m_{2}}$ with $\Psi_{ij}=\psi_{j}(\bm{y}_{i})$
while not converged do
  for each $k$:
    $\bm{b}^{(k)}=NN_{\theta}(\bm{a}^{(k)})$, $\hat{\bm{u}_{\bm{\theta}}}^{(k)}=\Psi\bm{b}^{(k)}$, $\bm{u}^{(k)}=\big(u^{(k)}(\bm{y}_{i})\big)_{i=1}^{n_{2}}$
  loss:
   $L(\bm{\theta})=\frac{1}{M}\sum_{k=1}^{M}\frac{\|\hat{\bm{u}_{\bm{\theta}}}^{(k)}-\bm{u}^{(k)}\|_{2}}{\|\bm{u}^{(k)}\|_{2}}$
  update: $\bm{\theta}\leftarrow\bm{\theta}-\eta\nabla_{\bm{\theta}}L(\bm{\theta})$
end while
Output: $N\!N_{\bm{\theta}}$
3 Basis Systems and Coefficient Estimation
The goal of operator learning is to approximate an underlying mapping $G:\mathcal{X}\mapsto\mathcal{Y},$ where both $\mathcal{X}$ and $\mathcal{Y}$ are infinite-dimensional function spaces. Directly learning such a map is intractable, so it is necessary to embed $\mathcal{X}$ and $\mathcal{Y}$ into finite-dimensional subspaces. In our proposed FB-C2CNet, this embedding is realized by choosing an input function system $\{\phi_{j}(\bm{x})\}_{j=1}^{m_{1}}$ and an output function system $\{\psi_{j}(\bm{y})\}_{j=1}^{m_{2}}$, which span the subspaces
$V_{\mathrm{in}}=\mathrm{span}\{\phi_{j}\}_{j=1}^{m_{1}}\quad\text{and}\quad V_{\mathrm{out}}=\mathrm{span}\{\psi_{j}\}_{j=1}^{m_{2}}.$
Each source function $f\in\mathcal{X}$ and target function $u=G(f)\in\mathcal{Y}$ can then be projected onto these subspaces as
$f(\bm{x})\approx\sum_{j=1}^{m_{1}}a_{j}\phi_{j}(\bm{x}),\qquad u(\bm{y})\approx\sum_{j=1}^{m_{2}}b_{j}\psi_{j}(\bm{y}),$
where $\mathbf{a}=[a_{1},\dots,a_{m_{1}}]\in\mathbb{R}^{m_{1}}$ and $\mathbf{b}=[b_{1},\dots,b_{m_{2}}]\in\mathbb{R}^{m_{2}}$ are the corresponding coefficient representations. Hence, the infinite-dimensional operator learning problem is transformed into learning a finite-dimensional coefficient mapping
$NN_{\bm{\theta}}:\mathbb{R}^{m_{1}}\to\mathbb{R}^{m_{2}},$
which approximates the coefficient-to-coefficient relationship induced by $G$. This is the core idea of our proposed Fixed-Basis Coefficient-to-Coefficient Operator Network (FB-C2CNet).
Therefore, in our proposed FB-C2CNet, two key aspects require particular attention: (i) the selection of the fixed basis function spaces, and (ii) the computation of the corresponding coefficients. In this section, we provide a detailed discussion of these two components.
3.1 Fixed Basis Systems
One of the key advantages of our proposed FB-C2CNet is that the input space $\mathcal{X}$ and the output space $\mathcal{Y}$ are directly represented using pre-specified basis functions. Since these basis functions (or the associated encoder) are fixed in advance, the network does not need to learn them during training. This design substantially lowers the training cost, as the learning process is reduced to estimating the coefficients with respect to the fixed bases, rather than simultaneously optimizing both the representation and the mapping. In this work, we primarily employ the random feature model and the finite element method as the basis function systems for constructing $V_{\text{in}}$ and $V_{\text{out}}$. In the following, we provide a brief introduction to these two basis function systems.
Random feature model
The random feature model (RFM) can be expressed as
$\phi(\bm{x};\bm{\theta})=\sum_{j=1}^{M}a_{j}\sigma(\bm{k}_{j}\cdot\bm{x}+b_{j}),$ (9)
where the inner parameters $\{\bm{k}_{j},b_{j}\}_{j=1}^{M}$ are randomly initialized, typically from a uniform distribution with $\bm{k}_{j}\sim U\left([-R_{m},R_{m}]^{d}\right)$ and $b_{j}\sim U\left([-R_{m},R_{m}]\right)$ [22]. The activation function $\sigma$ is often chosen as $\tanh(x)$ or $\sin(x)$. Importantly, only the outer coefficients $\{a_{j}\}_{j=1}^{M}$ are trainable, enabling the resulting optimization problem to be solved efficiently via standard linear algebra techniques such as least squares or QR decomposition. To further improve approximation, the partition of unity (PoU) method is often employed [22], leading to
$u(\bm{x})=\sum_{n=1}^{M_{p}}\omega_{n}(\bm{x})\sum_{j=1}^{J_{n}}a_{nj}\,\sigma(\bm{k}_{nj}\cdot\bm{x}+b_{nj}),$ (10)
where $\psi_{n}(\bm{x})$ are construction functions forming a PoU, and $\phi_{nj}(\bm{x})=\sigma(\bm{k}_{nj}\cdot\bm{x}+b_{nj})$ are random features. Here, $M_{p}$ denotes the number of partitions and $J_{n}$ the number of features in each partition.
In one dimension, $\psi_{n}$ can be either the characteristic function
$\omega_{n}^{a}(x)=\chi_{\{-1\leq\tilde{x}<1\}},$ (11)
or the smooth function
$\omega_{n}^{b}(x)=\begin{cases}\frac{1+\sin(2\pi\tilde{x})}{2},&-\tfrac{5}{4}\leq\tilde{x}<-\tfrac{3}{4},\\[6.0pt] 1,&-\tfrac{3}{4}\leq\tilde{x}<\tfrac{3}{4},\\[6.0pt] \frac{1-\sin(2\pi\tilde{x})}{2},&\tfrac{3}{4}\leq\tilde{x}<\tfrac{5}{4},\\[6.0pt] 0,&\text{otherwise},\end{cases}$ (12)
where $\tilde{x}=(x-x_{n})/r_{n}$ rescales the interval to [1,1][-1,1]. For $d$ dimensions, we set $\omega_{n}(\bm{x})=\prod_{k=1}^{d}\omega_{n}(x_{k}).$
A one-dimensional illustration is provided in Figure 5(a). This PoU-based RFM provides a flexible and efficient approximation framework for PDE solutions. Owing to its computational convenience, the random feature model has found applications across a broad range of problems [38, 39, 40]. Moreover, since it is a meshfree method, it can naturally handle complex computational domains without the need for mesh generation. In practice, it also exhibits advantages in high-dimensional problems, where its structure facilitates more tractable implementation and efficient approximation [41, 42, 43].
Finite element method
The finite element method (FEM) [44, 45] provides another natural choice of basis functions, particularly well-suited for approximating PDE solutions. Given a mesh $\mathcal{T}_{h}$ over the computational domain $\Omega\subset\mathbb{R}^{d}$, the FEM basis functions are constructed locally on each element. For example, in the case of linear ($P_{1}$) elements, the basis functions are the so-called hat functions, defined by
$\phi_{j}(\bm{x}_{i})=\delta_{ij},\quad 1\leq i,j\leq N_{h},$ (13)
where $\{\bm{x}_{i}\}_{i=1}^{N_{h}}$ are the nodal points of the mesh and $N_{h}$ is the number of degrees of freedom. The global FEM approximation of a function $u(\bm{x})$ can then be expressed as
$u_{h}(\bm{x})=\sum_{j=1}^{N_{h}}a_{j}\,\phi_{j}(\bm{x}),$ (14)
where the coefficients $\{a_{j}\}_{j=1}^{N_{h}}$ correspond to the nodal values of $u$. Higher-order elements ($P_{k}$, $k\geq 2$) can also be employed to improve approximation accuracy, in which case the basis functions are polynomials of degree $k$ defined on each element of the mesh. A one-dimensional illustration is provided in Figure 5(b). Due to their locality, conformity, and well-established approximation properties, FEM basis functions are particularly effective for incorporating boundary conditions and ensuring numerical stability. They have therefore become a standard tool for representing PDE solution spaces in scientific computing. Nevertheless, in high-dimensional settings they suffer from the curse of dimensionality, as the number of degrees of freedom grows exponentially with the dimension.
Other basis functions
In addition to FEM and random feature models, a variety of other basis function systems can be employed for constructing approximation spaces. Classical spectral methods, for instance, utilize global orthogonal polynomials such as Chebyshev or Legendre polynomials, which offer exponential convergence rates for smooth solutions [46, 47]. Alternatively, radial basis functions (RBFs), including Gaussian, multiquadric, and inverse multiquadric kernels, provide meshfree approximations that are particularly effective for scattered data and irregular domains [48, 49]. Wavelet bases have also been explored, offering multiresolution representations with good localization in both space and frequency [50]. Each of these choices comes with distinct advantages and trade-offs in terms of accuracy, computational cost, and adaptability to complex geometries, and they can be flexibly integrated into operator-learning frameworks depending on the problem setting.
Figure 3: *
(a) RFM basis.
Figure 4: *
(b) FEM basis.
Figure 5: Basis functions: (a) One-dimensional RFM example: $M_{p}=4$, with each partition containing $J_{n}=2$ random feature models, and random feature parameters sampled from the range $R_{m}=3$. (b) One-dimensional FEM example with $N_{h}=8$ nodal basis functions.
The choice of basis functions is problem-dependent and plays a crucial role in the accuracy and efficiency of the proposed framework. For simple one- and two-dimensional problems, finite element basis functions are often a natural and reliable choice, as they provide controllable accuracy and a solid theoretical foundation. For problems with complex geometries or irregular boundaries, finite element methods remain applicable, but constructing an appropriate mesh becomes highly non-trivial in practice, both in terms of the manual effort required and the associated computational cost. In such cases, mesh-free approaches based on random feature models offer a more flexible alternative. Moreover, for high-dimensional problems, the finite element method suffers from the curse of dimensionality, as the number of required basis functions grows exponentially with the spatial dimension. The random feature model, by contrast, avoids this issue and thus provides a more scalable and practical choice for high-dimensional operator learning. In the following numerical experiments, we present a series of cases comparing RFM-C2C and FEM-C2C, from which useful insights can be drawn regarding the choice of basis functions.
3.2 Coefficient representations and learnability
Once a fixed function space has been chosen, the next step is to determine the expansion coefficients $\{a_{j}\}_{j=1}^{m_{1}}$. At this stage, two fundamental questions naturally arise:
  • 1.
    What characterizes a desirable coefficient representation?
  • 2.
    How can the coefficients be computed efficiently and stably?
In the following, we address these questions in turn.
Coefficient Quality
We begin by examining the properties of coefficient representations that facilitate effective learning in the neural operator. The choice of coefficients is not merely a technical detail; it directly influences the trainability, convergence speed, and generalization ability of FB-C2CNet. For instance, poorly scaled or highly correlated coefficients may lead to ill-conditioned optimization landscapes, thereby making training unstable or inefficient. In contrast, well-structured coefficients that capture the essential variability of the target functions can significantly ease the learning burden of the network. In the following, we analyze what makes a coefficient representation “easy to learn,” and provide guidelines for constructing coefficient systems that are both expressive and numerically stable.
Suppose a set of basis functions $\{\phi_{j}(x)\}_{j=1}^{m_{1}}$ is chosen to span the input approximation space $V_{\mathrm{in}}$. Each training sample $f$ is projected onto this basis, yielding
$f(\mathbf{x})\approx\sum_{j=1}^{m_{1}}a_{j}\phi_{j}(\mathbf{x}),$
where $\bm{a}_{\mathrm{in}}=[a_{1},a_{2},\dots,a_{m_{1}}]^{\top}$ denotes the coefficient vector.Given a training dataset consisting of $M$ input functions $\{f^{(k)}(\mathbf{x})\}_{k=1}^{M}$, we obtain their corresponding coefficient representations $\{\bm{a}^{(k)}\}_{k=1}^{M}$, where
$\bm{a}^{(k)}=[a_{1}^{(k)},a_{2}^{(k)},\dots,a_{m_{1}}^{(k)}]^{\top}\in\mathbb{R}^{m_{1}\times 1}.$
Collecting all coefficient vectors into a matrix form gives
$A=\begin{bmatrix}a_{1}^{(1)}&a_{2}^{(1)}&\cdots&a_{m_{1}}^{(1)}\\[3.0pt] a_{1}^{(2)}&a_{2}^{(2)}&\cdots&a_{m_{1}}^{(2)}\\[3.0pt] \vdots&\vdots&\ddots&\vdots\\[3.0pt] a_{1}^{(M)}&a_{2}^{(M)}&\cdots&a_{m_{1}}^{(M)}\end{bmatrix}\in\mathbb{R}^{M\times m_{1}},$
Each row of $A$ corresponds to a projected training function in the coefficient space of $V_{\mathrm{in}}$.
A useful diagnostic for assessing the learnability of such coefficient representations is the effective rank of the matrix $A$.
Definition 1 (Effective rank [51]).
For a matrix $\bm{A}\in\mathbb{R}^{m\times n}$ $(m>n)$ with singular values $\sigma_{1}\geq\sigma_{2}\geq\ldots\geq\sigma_{n}$, the truncation entropy of $\bm{A}$ is defined as:
$\operatorname{erank}(\bm{A})=\exp\left\{-\sum_{k=1}^{n}p_{k}\log p_{k},\right\},$ (15)
where $p_{k}=\frac{\sigma_{k}}{\|\bm{\sigma}\|_{1}}$, for $k=1,2,\ldots,n$, with $\bm{\sigma}=[\sigma_{1},\sigma_{2},\dots,\sigma_{n}]^{\top}$.
The effective rank can measure the intrinsic dimensionality of the coefficient distribution. Intuitively, a larger effective rank indicates that the coefficients span a richer and more balanced subspace of $\mathbb{R}^{m_{1}}$, which eases optimization and improves generalization. In contrast, when the effective rank is small, the coefficients concentrate in a low-dimensional subspace, leading to poor conditioning, slow convergence, and increased risk of overfitting.
In addition to the effective rank, the distribution of variations in the coefficient vector $\mathbf{a}$ associated with the basis functions $\{\phi_{j}(\mathbf{x})\}_{j=1}^{m_{1}}$ also plays an important role [52]. When the coefficients exhibit large fluctuations along only a few dominant basis directions, the learned model tends to overfit these components and generalizes poorly to unseen inputs. In contrast, when the variations in $\mathbf{a}$ are more evenly distributed among the basis functions, the representation becomes smoother and more balanced, leading to improved generalization performance.
Consequently, in the process of computing coefficients, our objective is to obtain a coefficient matrix that possesses both high effective rank and balanced variance distribution. Such properties not only enhance the expressiveness of the representation but also improve the trainability and generalization ability of FB-C2CNet.
Coefficient Computation
After fixing the basis functions, we proceed to the computation of the associated coefficients. Since the basis functions $\{\phi_{j}(\mathbf{x})\}_{j=1}^{m_{1}}$ are fixed and known, the Equation 4 can be solved directly in matrix form. Define the design matrix $\Phi\in\mathbb{R}^{n_{1}\times m_{1}}$ as
$\Phi_{ij}=\phi_{j}(\mathbf{x}_{i}),\quad 1\leq i\leq n_{1},\;1\leq j\leq m_{1},$ (16)
and let the input function values be $\mathbf{f}=\big[f(x_{1}),f(x_{2}),\dots,f(x_{n_{1}})\big]^{\top}\in\mathbb{R}^{n_{1}\times 1}.$ Then the coefficient vector $\mathbf{a}\in\mathbb{R}^{m_{1}\times 1}$ is determined from
$\Phi\mathbf{a}=\mathbf{f}.$ (17)
In our framework, the goal is to reduce the input dimensionality, and thus we typically encounter the case $m_{1}\ll n_{1}$, where the number of basis functions is much smaller than the number of sampling points. In this case, we compute $\mathbf{a}$ by solving the regularized least-squares problem
$\mathbf{a}^{\ast}=\arg\min_{\mathbf{a}\in\mathbb{R}^{m_{1}}}\|\Phi\mathbf{a}-\mathbf{f}\|_{2}^{2}+\lambda\|\mathbf{a}\|_{2}^{2},$ (18)
which admits the closed-form solution
$\mathbf{a}^{\ast}=(\Phi^{\top}\Phi+\lambda I)^{-1}\Phi^{\top}\mathbf{f}.$ (19)
The regularization parameter $\lambda$ controls the smoothness of the recovered coefficients: a larger $\lambda$ enforces stronger regularity and suppresses small-scale variations in $\mathbf{a}$.
Alternatively, the same regularization effect can be realized through the singular value decomposition (SVD). Let $\Phi=U\Sigma V^{\top}$ be the SVD of $\Phi$, with singular values $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{m_{1}}$. A truncated SVD (or cut-off) approach constructs the pseudo-inverse by discarding singular values below a threshold $\sigma_{\text{cut}}$, yielding
$\mathbf{a}^{\ast}=V\Sigma^{\dagger}U^{\top}\mathbf{f},$ (20)
where $\Sigma^{\dagger}=\mathrm{diag}(\sigma_{1}^{-1},\dots,\sigma_{r}^{-1},0,\dots,0)$ and $r$ is the number of retained singular modes. The cut-off parameter $\sigma_{\text{cut}}$ plays a role analogous to the regularization parameter $\lambda$: a larger cut-off (corresponding to stronger regularization) removes more high-frequency components, leading to smoother but less detailed coefficient representations.
From a spectral perspective, increasing the cut-off threshold effectively reduces the contribution of small singular values, resulting in a higher effective rank of the matrix $\Phi$ and reduced variations in the coefficient vector $\mathbf{a}$. Hence, the choice of $\sigma_{\text{cut}}$ (or equivalently, $\lambda$) directly controls the balance between stability, smoothness, and the expressive richness of the learned representation.
We observe that when the random feature model (RFM) is used as the basis, the cut-off strategy has a particularly significant impact on the resulting coefficients. This sensitivity arises because random features often lead to highly correlated basis functions, causing the singular value spectrum of $\Phi$ to decay rapidly [13]. As a result, the choice of truncation threshold strongly influences the quality of the coefficient representation.
Example
1D Darcy flow as an example to illustrate our finding.
To gain further intuition, we conduct a simple experiment to illustrate the effect of the SVD cut-off on the effective rank of the coefficient matrix $A$ in the RFM setting. Specifically, we consider the one-dimensional Darcy flow problem introduced in [27]. For this test, we set the number of random features in both the input and output basis systems to $m_{1}=m_{2}=128$, and discretize the domain with $1000$ uniformly spaced grid points. Figure 9 shows the resulting coefficient distributions under different cut-off thresholds. We observe that as the cut-off increases, the coefficients become more concentrated, the variance across different modes decreases, and the effective rank correspondingly grows. This confirms that in the RFM setting, a larger cut-off can paradoxically lead to coefficient matrices with higher effective rank, owing to the rapid decay of the singular value spectrum.
Figure 6: *
Figure 7: *
Figure 8: *
Figure 9: Coefficient distributions. Illustration of the coefficient distributions obtained under different SVD cut-off thresholds in the RFM setting. As the cut-off increases, the coefficients become more concentrated, the variance across modes decreases, and the effective rank correspondingly grows.
As shown in Figure 10, although the training curves are nearly identical across different cut-off thresholds, the generalization performance exhibits a clear dependence on the effective rank of the coefficient matrix. When the cut-off is small, the resulting coefficients yield a low effective rank, indicating that the representation is concentrated in a limited number of directions. Such low-rank structure restricts expressiveness and leads to poor generalization. By contrast, larger cut-off thresholds increase the effective rank, producing coefficient representations that span a richer subspace of the function space. This higher effective rank is consistently associated with lower test error, demonstrating that effective rank is a key indicator of generalization performance in FB-C2CNet.
Figure 10: Train and test curve. Illustration of the training and testing performance of FB-C2CNet when using coefficients computed with different SVD cut-off thresholds as inputs. The results show that the choice of cut-off significantly influences the coefficient representation, and consequently the training dynamics and generalization ability of the network.
3.3 Output-Space-Induced Error Analysis
In the following, we demonstrate that the ultimate accuracy of the proposed method is fundamentally constrained by the approximation capability of the output space $V_{\mathrm{out}}$. Let $u(\bm{y})$ denote the ground-truth solution of the target operator $G:\mathcal{X}\to\mathcal{Y}$, i.e., $u(\bm{y})\in\mathcal{Y}$. Once the output function system is fixed as $V_{\mathrm{out}}=\mathrm{span}\{\psi_{j}(\bm{y})\}_{j=1}^{m_{2}},$ any $u(\bm{y})\in\mathcal{Y}$ admits an approximation
$\hat{u}(\bm{y})=\sum_{j=1}^{m_{2}}b_{j}\psi_{j}(\bm{y}),\qquad\hat{u}\in V_{\mathrm{out}},$ (21)
where $\bm{b}=[b_{1},b_{2},\dots,b_{m_{2}}]$ denotes the output coefficient vector. The FB-C2CNet is trained to learn the coefficient-to-coefficient mapping
$NN_{\theta}:\bm{a}\in\mathbb{R}^{m_{1}}\to\bm{b}\in\mathbb{R}^{m_{2}},$
so that the reconstructed prediction reads
$\hat{u}_{\bm{\theta}}(\bm{y})=NN_{\bm{\theta}}(\bm{a})\cdot\Psi,\qquad\Psi=[\psi_{1}(\bm{y}),\psi_{2}(\bm{y}),\dots,\psi_{m_{2}}(\bm{y})]\in\mathbb{R}^{n_{2}\times m_{2}}.$ (22)
We then consider the following triangle inequality:
$\|u-\hat{u}\|_{L^{2}}\;\leq\;\underbrace{\|\hat{u}-\hat{u}_{\theta}\|_{L^{2}}}_{\text{learnable operator error}}+\underbrace{\|u-\hat{u}_{\theta}\|_{L^{2}}}_{\text{error for the FB-C2CNet}}.$ (23)
Here, the first term $\|\hat{u}-\hat{u}_{\theta}\|_{L^{2}}$ measures the learnable discrepancy introduced by the neural mapping $NN_{\theta}$, indicating how well the network reproduces the coefficient-to-coefficient relation. The second term $\|u-\hat{u}_{\theta}\|_{L^{2}}$ represents the overall approximation error of the FB-C2CNet with respect to the reference solution $u$, which includes both the learnable error and the representation error associated with $V_{\mathrm{out}}$.
We take the lower bound of the left-hand side and denote it as the intrinsic limitation imposed by the output space $V_{\mathrm{out}}$, which is characterized by the projection error:
$\varepsilon_{\mathrm{proj}}:=\inf_{\mathbf{b}\in\mathbb{R}^{m_{2}}}\bigg\|u(\bm{y})-\sum_{j=1}^{m_{2}}b_{j}\psi_{j}(\bm{y})\bigg\|_{L^{2}},$ (24)
which satisfies $\varepsilon_{\mathrm{proj}}\leq\|u-\hat{u}\|_{L^{2}}$ for any choice of coefficients. This quantity can be regarded as a measure of the approximation capability of $V_{\mathrm{out}}$ to represent functions in $\mathcal{Y}$. In practice, the coefficients $\bm{b}$ are obtained via a least-squares projection of the target function onto the subspace $V_{\mathrm{out}}$. Combining with the above triangle inequality yields
$\varepsilon_{\mathrm{proj}}\;\leq\;\|\hat{u}-\hat{u}_{\theta}\|_{L^{2}}+\|u-\hat{u}_{\theta}\|_{L^{2}}.$ (25)
Therefore, under the assumption that the FB-C2CNet is sufficiently well trained such that the learnable mapping $NN_{\theta}$ accurately captures the underlying operator $G$, the total approximation error is dominated by, and asymptotically bounded by, the projection error associated with the output space $V_{\mathrm{out}}$, i.e.,
$\varepsilon_{\mathrm{proj}}\;\leq\;\|u-\hat{u}_{\theta}\|_{L^{2}}.$ (26)
In the following numerical experiments, we will further verify that this theoretical prediction is consistent with the empirical results.
4 Numerical Experiments
In this section, we present a series of experiments across a broad spectrum of PDE settings to demonstrate the effectiveness of our method: (i) the Darcy flow problem on regular domains in both one and two dimensions, (ii) the Poisson equation on a two-dimensional complex domain, (iii) 2D Darcy flow on a complex domain, (iv) elastic plate problems, and (v) high-dimensional Poisson equations on regular domains. These examples encompass diverse dimensionalities, domain geometries, and operator learning scenarios, following the setups in [35, 36].
It is worth noting that both the elastic plate and 2D Darcy flow problems in complex domains belong to the category of multi-input/output operator learning. In the elastic plate case, a single input function (the load distribution) is mapped to two output functions representing displacement components, while in the 2D Darcy flow case, two input functions (the permeability field and source term) are mapped to a single output pressure field. In these settings, we generalize the proposed framework to handle multi-component mappings and compare two encoding strategies: a scalar-valued representation formed by concatenating the coefficients of multiple inputs, and a vector-valued representation that employs shared vector-valued basis functions to jointly encode all components.
Unless otherwise stated, the neural operator is implemented as a fully connected network with architecture $m_{1},512,m_{2}]$, where $m_{1}$ represents the number of basis functions used to approximate the input function space, and $m_{2}$ denotes the number of basis functions used to approximate the output function space. All models are trained with the Adam optimizer, employing a learning rate schedule that decays from $10^{-3}$ to $10^{-6}$ (as illustrated in Figure 58). All experiments are performed on a single NVIDIA RTX 4090 GPU. We utilize the Gmsh Python library for mesh generation. The resulting mesh is then used to construct the FEM basis functions. Additional implementation details and hyperparameter settings are provided in the Appendix.
4.1 1D Darcy Flow
In this example, we aim to learn the nonlinear Darcy operator for a one-dimensional system. A variant of the nonlinear 1D Darcy equation is given by
$\frac{d}{dx}\left(a(u)\frac{du}{dx}\right)$ $=f(x),\quad x\in[0,1],$ (27)
$u(0)=u(1)$ $=0,$
where the solution-dependent permeability is $a(u(x))=0.2+u^{2}(x)$ and the input term $f$ is modeled as a Gaussian random field, $f\sim\mathcal{GP}(0,k(x,x^{\prime}))$, with covariance kernel $k(x,x^{\prime})=\sigma^{2}\exp\left(-\frac{|x-x^{\prime}|^{2}}{\ell^{2}}\right)$, $\ell=0.04$, and $\sigma=1.0$. The dataset is from [36, 37]. Our objective is to learn the map:
$G:f(x)\mapsto u(x).$
The training set consists of 500 samples and the testing set consists of 200 samples. The spatial domain $[0,1]$ is discretized uniformly with 2000 grid points.
C2C vs. P2C: Efficiency and Accuracy. In this example, we compare the coefficient-to-coefficient (C2C) and point-to-coefficient (P2C) approaches using basis functions from the finite element method and the random feature model. For the P2C setting, the network architecture and training procedure remain identical to those of the C2C framework, except that the neural network takes pointwise function values as inputs instead of coefficients. This design highlights the effectiveness of using coefficient representations as neural network inputs. As shown in Figure 13, with 2000 spatial points, the P2C method requires a high-dimensional input (dimension 2000), resulting in a larger number of network parameters and more difficult training, which leads to slower loss decay and lower accuracy. In contrast, the proposed C2C method uses far fewer parameters and achieves higher accuracy, highlighting its efficiency.
Figure 11: *
(a)RFM basis.
Figure 12: *
(b) FEM basis.
Figure 13: 1D Darcy Flow: Training curve comparison between the C2C and P2C methods with different basis functions. The number of grid points is set to $n_{1}=n_{2}=2000$, and the number of basis functions is $m_{1}=128$ and $m_{2}=32$.
Generalization through different resolutions. We train both RFM-C2C and FEM-C2C on data generated with $n_{\mathrm{train}}=100$ grid points, and then evaluate their performance on test datasets with $n=40$, $500$, $1000$, and $2000$ grid points, respectively. This experiment examines the ability of the learned operators to generalize across different spatial resolutions. The results, summarized in Fig. 14, show that the proposed RFM-C2C maintains stable accuracy across unseen resolutions.
Figure 14: 1D Darcy Flow: FB-C2CNet is trained on a uniform $100$-point grid and tested on different resolutions ($n=40,500,1000,2000$). The network exhibits strong resolution invariance with consistent test accuracy across all cases.
4.2 2D Darcy Flow in regular domain
In this example, we consider a two-dimension Darcy flow equation in the square domain $\Omega=[0,1]^{2}$. It is defined as
$-\nabla\cdot\left(a(x,y)\nabla u(x,y)\right)$ $=f(x,y),\quad(x,y)\in\Omega,$ (28)
$u(x,y)$ $=0,\quad(x,y)\in\partial\Omega,$
where $a(x,y)$ is the diffusion coefficient and $f(x,y)$ is the forcing function. Here we set $f(x,y)\equiv 1$. Our objective is to learn the map:
$G:a(x,y)\mapsto u(x,y).$
The diffusion coefficient $a(x,y)$ is sampled from random field $exp(\psi)$, where $\psi=\mathcal{N}(0,(-\Delta+9I)^{-2})$. The dataset is from [9, 35]. The training set consists of 2400 samples and the testing set consists of 600 samples. The spatial domain $0,1]\times[0,1]$ is discretized uniformly with $141\times 141$ grid points.
C2C vs. P2C: Efficiency and Accuracy. In this example, similar to the 1D Darcy flow case, we compare the C2C and P2C approaches with different choices of basis functions. As shown in Figure 17, the results indicate that the C2C method achieves lower training error, which in turn leads to higher accuracy.
Figure 15: *
(a) RFM basis.
Figure 16: *
(b) FEM basis.
Figure 17: 1D Darcy Flow: Training curve comparison between the C2C and P2C methods with different basis functions. The number of grid points is set to $n_{1}=n_{2}=141\times 141$, and the number of basis functions is $m_{1}=1024$ and $m_{2}=2048$.
4.3 Poisson equation in complex domain
In this example, we consider a two-dimensional Poisson equation defined as
$\Delta u(x)$ $=f(x),\quad x\in\Omega,$ (29)
$u(x)$ $=0,\quad x\in\partial\Omega,$
where the computational domain $\Omega$ is a rectangle with three interior holes of different shapes. The objective is to learn the solution operator
$G:f(x)\mapsto u(x).$
The forcing term $f(x)$ is sampled from the function space
$f(x)=\sum_{i=0}^{4}\sum_{j=0}^{4}\alpha_{ij}\sin\!\left([i\pi,j\pi]\cdot x^{\top}\right),$
where the coefficients $\alpha_{ij}$ are independently drawn from the uniform distribution on [1,1][-1,1]. The reference solutions are generated using the finite element method (FEM) on a mesh with 1780 nodes as shown in Figure 27 (a).
In this experiment, we employ both FEM-C2C and RFM-C2C. The finite element results are shown in Figure 20 and Figure 24, which illustrate the worst-case comparison in the test dataset together with the corresponding pointwise error. Similarly, the RFM results are presented in Figure 27 and Figure 31, also demonstrating the worst-case comparison in the test dataset and the associated pointwise error.
Relation between Basis Approximation and FB-C2C Error. In this experiment, we aim to verify that the accuracy of the proposed FB-C2C method is influenced by the approximation capability of the output basis system. As illustrated in Figure 20(b) and Figure 27(b), the red dashed line represents the approximation error of the output basis system to the output function space, while the blue solid line represents the accuracy of the FB-C2C method in solving the problem. It can be observed that the blue curve always lies above the red dashed curve, indicating that the accuracy of FB-C2C is bounded by the approximation accuracy of the output basis system (i.e., the error cannot be smaller than the approximation error). In this experiment, increasing the number of output basis functions $m_{2}$ generally improves the approximation capability of the output basis system. Consequently, as $m_{2}$ increases, both the approximation error and the FB-C2C error decrease. The fact that the two curves nearly coincide further demonstrates that FB-C2C is sufficiently trained in this setting.
Figure 18: *
(a)FEM basis for output function space.
Figure 19: *
(b) FEM basis.
Figure 20: Poisson equation in complex domain: FEM results. (a) Finite element mesh of the output basis system. (b) The FEM-C2C error (blue) is bounded below by the approximation error of the output basis system (red), and both decrease as the number of output bases $m_{2}$ increases.
Figure 21: *
Figure 22: *
Figure 23: *
Figure 24: Poisson equation in complex domain: Worst-case sample from the complex domain Poisson dataset solved by FEM-C2C. (a) Pointwise comparison between the prediction and the ground truth. (b) Exact solution. (c) Predicted solution.
Figure 25: *
(a) Sample grid points for the functions.
Figure 26: *
(b) RFM basis.
Figure 27: Poisson equation in complex domain: RFM results. (a) Distribution of sampled points. (b) The RFM-C2C error (blue) is bounded below by the approximation error of the output basis system (red), and both decrease as the number of output bases $m_{2}$ increases.
Figure 28: *
Figure 29: *
Figure 30: *
Figure 31: Poisson equation in complex domain: Worst-case sample from the complex domain Poisson dataset solved by RFM-C2C. (a) Pointwise comparison between the prediction and the ground truth. (b) Exact solution. (c) Predicted solution.
4.4 Elastic plate
In this example, we demonstrate the capability of the proposed method to approximate solution operators for problems with coupled multi-field outputs. This investigation concerns a thin rectangular plate subjected to in-plane loading, described by the two-dimensional plane stress elasticity equations under the assumption of negligible body forces. The computational domain is defined as $\Omega=[0,1]\times[0,1]$ with a thin rectangular region removed from the interior, representing a cut-out in the plate. The governing equations are
$\begin{array}[]{cc}\nabla\cdot\mathbf{\sigma}+\mathbf{b}(\mathbf{x})=0,\quad\mathbf{x}=(x,y),&\\ (u,v)=0,\quad\forall\,x=0,&\\ \sigma\cdot\mathbf{n}=f(\mathbf{x}),\quad\forall\,x=1,\end{array}$ (30)
where $\sigma$ denotes the Cauchy stress tensor, $\mathbf{b}(\mathbf{x})=0$ is the body force, $\mathbf{n}$ is the unit outward normal vector, $f(\mathbf{x})$ is the boundary load, and $u(\mathbf{x})$ and $v(\mathbf{x})$ represent the displacements in the $x$- and $y$-directions, respectively. The material parameters are the Young’s modulus $E=300\times 10^{5}$ and Poisson’s ratio $\nu=0.3$. Under plane stress conditions, the constitutive relation between stress and displacement gradients is given by
$\begin{bmatrix}\sigma_{xx}\\ \sigma_{yy}\\ \tau_{xy}\end{bmatrix}=\frac{E}{1-\nu^{2}}\begin{bmatrix}1&\nu&0\\ \nu&1&0\\ 0&0&\frac{1-\nu}{2}\end{bmatrix}\begin{bmatrix}\displaystyle\frac{\partial u}{\partial x}\\[8.61108pt] \displaystyle\frac{\partial v}{\partial y}\\[8.61108pt] \displaystyle\frac{\partial u}{\partial y}+\frac{\partial v}{\partial x}\end{bmatrix}$. (31)
The boundary loading $f(1,y)$ applied to the right edge of the plate is modeled as a Gaussian random field. Our objective is to learn the map:
$G:f(1,y)\mapsto[u(\mathbf{x}),v(\mathbf{x})].$
The dataset for this problem is adopted from [36]. The input grid points are uniformly sampled over the domain with $101$ points, while the test grid contains $1048$ points, which is shown in Figure 35 (a). In total, $1850$ samples are used for training and $100$ samples are reserved for testing.
We now examine an operator with a multi-function output (specifically, two functions), whereas previous cases involved single-output operators. The following section details the two approaches pursued for this scenario: the scalar-valued and vector-valued basis methods.
Scalar-Valued basis functions for Multi-Output Function Approximation. We use scalar-valued basis functions $\{\psi_{i}(\mathbf{x})\}_{i=1}^{m_{2}}$ to approximate $u(\mathbf{x})$ and $v(\mathbf{x})$, i.e.,
$u(\mathbf{x})\approx\sum_{j=1}^{m_{2}}b_{j}^{u}\,{\psi}_{j}(\mathbf{x}),\quad v(\mathbf{x})\approx\sum_{j=1}^{m_{2}}b_{j}^{v}\psi_{j}(\mathbf{x}),$ (32)
where the corresponding output coefficients $\mathbf{b}=[b_{1}^{u},b_{2}^{u},..,b_{m_{2}}^{u},b_{1}^{v},b_{2}^{v},...,b_{m_{2}}^{v}]\in\mathbb{R}^{m_{2}\times 2}$ associated with the scalar-valued basis space $V_{\text{out}}=\text{span}\{\psi_{i}(\mathbf{x})\}_{i=1}^{m_{2}}$.
Vector-Valued basis functions for Multi-Output Function Approximation. In this example, we denote vector-valued basis space
$V_{\text{out}}=\text{span}\{\bm{\psi}_{i}(\mathbf{x})\}_{i=1}^{m_{2}}=\text{span}\{[\psi_{i}^{u}(\mathbf{x}),\psi_{i}^{v}(\mathbf{x})]\}_{i=1}^{m_{2}},$
which are employed to approximate the multi-output function space $\mathcal{Y}$. A given output multiple functions $\bm{u}(x)=[u(x),v(x)]\in\mathcal{Y}$ is represented in terms of the vector-valued basis functions, i.e.,
$\bm{u}(\mathbf{x})=[u(\mathbf{x}),v(\mathbf{x})]\approx\sum_{j=1}^{m_{2}}b_{j}\,\bm{\psi}_{j}(\mathbf{x})=\sum_{j=1}^{m_{2}}b_{j}\,[\psi_{j}^{u}(\mathbf{x}),\psi_{j}^{v}(\mathbf{x})],$ (33)
where $\mathbf{b}=[b_{1},b_{2},...,b_{m_{2}}]\in\mathbb{R}^{m_{2}}$ are the expansion coefficients associated with multiple output functions $\bm{u}=[u,v]$.
Scalar-Valued vs. Vector-Valued Basis Functions: Efficiency and Accuracy. Both methods use $m_{1}=64$ input basis functions, a count precisely set in one dimension. Their output dimensions differ: in the 2D scalar-valued case, RFM attains $m_{2}=512$ while FEM yields $m_{2}=540$, as the latter cannot precisely mesh 512 triangles; in the vector-valued case, randomness generates a different basis, but these $m_{2}$ values are preserved. Two different meshes for the FEM basis generation are shown in Figure 35 (b) and Figure 35 (c).
As the vector-valued basis requires only half the output coefficients in this example, it contains significantly fewer trainable parameters, resulting in faster convergence and enhanced training efficiency. The relative error curves in Figure 38 (RFM) show similar accuracy for scalar-valued and vector-valued bases. Conversely, Figure 41 (FEM) reveals a significant accuracy gap, with the vector-valued basis performing much worse than the scalar-valued one. Therefore, the vector-valued basis provides superior training efficiency over the scalar-valued basis and, as realized in the RFM-C2C method, markedly better accuracy than the vector-valued FEM-C2C method.
Figure 32: *
(a) Sample grid points.
Figure 33: *
(b) $\psi$ $(\psi^{v})$ FEM basis mesh.
Figure 34: *
(c) $\psi^{u}$ FEM basis mesh.
Figure 35: Elastic plate: Grid points and two FEM basis meshes for the output functions.
Figure 36: *
(a) Scalar-Valued RFM basis.
Figure 37: *
(b) Vector-Valued RFM basis.
Figure 38: Training curve comparison between the scalar-valued and vector-valued RFM basis. The number of grid points is set to $n_{1}=101$ and $n_{2}=1048$, and the number of basis functions is $m_{1}=64$ and $m_{2}=512$.
Figure 39: *
(a) Scalar-Valued FEM basis.
Figure 40: *
(b) Vector-Valued FEM basis.
Figure 41: Training curve comparison between the scalar-valued and vector-valued FEM basis. The number of grid points is set to $n_{1}=101$ and $n_{2}=1048$, and the number of basis functions is $m_{1}=64$ and $m_{2}=540$.
4.5 2D Darcy Flow in complex domain
We investigate Darcy flow in a two-dimensional L-shaped region to validate our approach in handling complex domain configurations. A distinctive feature of this example is that it involves a mapping from two functions to a single function, thereby highlighting the broad applicability of our method. The problem is formulated as
[…]
$u(\mathbf{x})$ $=g(\mathbf{x}),\quad\mathbf{x}\in\partial\Omega,$
where $k(\mathbf{x})$, $u(\mathbf{x})$, and $f(\mathbf{x})$ denote the permeability field, hydraulic head, and source term, respectively, all exhibiting spatial variability. The dataset is from [36]. Its training set consists of 51000 samples and the testing set consists of 7000 samples. Figure 44 shows the discretization of input and output functions over sampling grids of 736 and 450 points, respectively. In this case we aim to approximate the map:
$G:[k(\mathbf{x}),f(\mathbf{x})]\mapsto u(\mathbf{x}).$
This example departs from the preceding cases by considering an operator that maps two input functions to a single output. To enhance both training efficiency and accuracy, we leverage vector-valued basis functions, specifically employing the RFM type motivated by our earlier findings.
Vector-Valued Basis Functions for Multi-Input Function Approximation. For the approximation of a single function, as considered in previous numerical experiments, we may use a linear space spanned by a set of scalar-valued basis functions. However, for the joint approximation of a family of functions consisting of multiple components, it is therefore better to work within a vector space spanned by vector-valued basis functions, as it leads to superior training efficiency.
In this example, we define the input basis system as follows.
$V_{\text{in}}=\text{span}\{\bm{\phi}_{i}(\mathbf{x})\}_{i=1}^{m_{1}}=\text{span}\{[\phi_{i}^{f}(\mathbf{x}),\phi_{i}^{k}(\mathbf{x})]\}_{i=1}^{m_{1}},$
which is used to approximate the input function space $\mathcal{X}$. A given input pair of functions $\mathbf{f}(\mathbf{x})=[f(\mathbf{x}),k(\mathbf{x})]\in\mathcal{X}$ can then be represented in terms of these vector-valued basis functions:
$\mathbf{f}(\mathbf{x})=[f(\mathbf{x}),k(\mathbf{x})]\approx\sum_{j=1}^{m_{1}}a_{j}\,\bm{\phi}_{j}(\mathbf{x})=\sum_{j=1}^{m_{1}}a_{j}\,[\phi_{j}^{f}(\mathbf{x}),\phi_{j}^{k}(\mathbf{x})],$ (35)
where $\{a_{j}\}_{j=1}^{m_{1}}$ are the corresponding expansion coefficients.
The coefficients $\{a_{j}\}$ are obtained by solving the following vector-valued regularized least-squares problem:
$\min_{\mathbf{a}\in\mathbb{R}^{m_{1}}}\;\frac{1}{n}\sum_{i=1}^{n}\left\|\mathbf{f}(\mathbf{x}_{i})-\sum_{j=1}^{m_{1}}a_{j}\,\bm{\phi}_{j}(\mathbf{x}_{i})\right\|_{2}^{2}+\lambda\|\mathbf{a}\|_{2}^{2},$ (36)
where $\{\mathbf{x}_{i}\}_{i=1}^{n}$ are the sampling points in the domain, and the second term provides $L^{2}$-regularization to stabilize the coefficient estimation.
Generalization Performance of RFM-C2C vs. RFM-P2C. In this example, we compare the generalization error of RFM-C2C and RFM-P2C methods by examining the effect of varying the number of training samples on the test results. The generalization capabilities of both methods under fixed RFM basis spaces are evaluated using training sets of 1000 samples, 5100 samples, and the complete dataset. As shown in Figure 48, the C2C method consistently achieves superior generalization performance and higher accuracy compared to the P2C method across all three training sets. This advantage becomes increasingly evident as the amount of training data decreases, demonstrating the robustness of the C2C method in data-scarce scenarios. Furthermore, the worst-case relative error from the L-shaped Darcy flow test set, obtained with full training data using the C2C method, is illustrated in Figure 48.
Figure 42: *
(a) Distribution of input function sample points.
Figure 43: *
(b) Distribution of output function sample points.
Figure 44: Darcy flow in L-shaped domain: Sample grid points for functions.
Figure 45: *
(a) 1000 training data
Figure 46: *
(b) 5100 training data
Figure 47: *
(c) 51000 training data
Figure 48: Training curve comparison between the RFM-C2C and RFM-P2C methods in three training datasets of varying scales. All three test datasets are identically sized, each comprising 7000 instances. The number of grid points is set to $n_{1}=736\times 2$ and $n_{2}=450$, and the number of RFM basis functions is $m_{1}=1024$ and $m_{2}=1024$.
Figure 49: *
Figure 50: *
Figure 51: *
Figure 52: Darcy flow in L-shape domain: Worst-case sample from the L-shape Darcy flow dataset with 51000 training data. (a) Pointwise comparison between the prediction and the ground truth. (b) Exact solution. (c) Predicted solution.
4.6 High Dimension Poisson Equation
In this example, we consider operator learning for high-dimensional functions. In this setting, if the finite element method is employed to construct the basis functions, one inevitably encounters the curse of dimensionality, as the number of basis functions required for an accurate approximation grows exponentially with the spatial dimension. Therefore, in such cases, adopting the RFM-C2C framework provides a more natural and efficient alternative. Also, in such cases, the number of grid points typically grows exponentially with the dimension, making conventional point-to-coefficient (P2C) frameworks impractical. By employing random feature models (RFM) as the basis functions, our approach effectively alleviates this issue and reduces the complexity of the neural operator network. To illustrate the capability of our method in addressing high-dimensional problems, we present the following numerical example on the high-dimensional Poisson equation:
$-\Delta u=$ $f\quad\text{in }\Omega,$ (37)
$u=\alpha\sum_{i=1}^{d}\sin\!\left(\tfrac{\pi}{2}x_{i}\right)$ $\quad\text{on }\partial\Omega,$
where $\Omega=[-1,1]^{d}$ and $f(\mathbf{x})=\alpha\frac{\pi^{2}}{4}\sum_{i=1}^{d}\sin\left(\frac{\pi}{2}x_{i}\right)$, which admits the exact solution
$u(\mathbf{x})=\alpha\sum_{i=1}^{d}\sin\left(\frac{\pi}{2}x_{i}\right).$
Here the objective is to learn the map
$G:f(\bm{x})\mapsto u(\bm{x}).$
To construct the training dataset, we uniformly sample $31^{d}$ points from the domain $1,1]^{d}$ and choose $8$ random values of $a\in[-1,1]$. The solutions are then evaluated for different values of $a$. In particular, we designate as Generation 1 (Gen1) the case of $2$ test values with $a\in[-1,1]$ that are not included in the training set. Furthermore, we perform extrapolation tests on $10$ test values of $a\in[1,2]$, which lie outside the training range and are designated as Generation 2 (Gen2). When training the RFM to obtain the coefficients, we employ the stochastic gradient descent (SGD) method. At each training step, we use a batch size of $1e5$, i.e., we randomly select $10^{5}$ points from the total $31^{10}$ grid points to compute the RFM coefficients.
In this example, we adopt the random feature model as the basis functions and use the same basis for both the input and output spaces, i.e., $m_{1}=m_{2}$. We compare two settings with $m_{1}=m_{2}=128$ and $m_{1}=m_{2}=8192$, respectively. The results are summarized in Table 1. From the results in Table 1, we observe that our RFM-based C2C method is able to effectively handle high-dimensional problems while achieving strong generalization. Using a smaller number of basis functions can accelerate training, though at the expense of a slight loss in accuracy. The training and testing curves are presented in Figure 53, while the worst-case sample is illustrated in Figure 57.
$m_{1}=m_{2}=128$ $m_{1}=m_{2}=8192$
Dimension $d$ Gen1 Gen2 Training Time Gen1 Gen2 Trianing Time
10 $1.69\mathrm{e\text{-}02}$ $3.00\mathrm{e\text{-}02}$ 1208.5s $6.48\mathrm{e\text{-}03}$ $1.27\mathrm{e\text{-}02}$ 43403.6s
Table 1: High-dimensional Poisson: Relative $L^{2}$ errors between the predicted and exact solutions $u_{i}(\mathbf{x})$ for different numbers of basis functions. The training time refers to the time required to train the neural operator.
Figure 53: High-Dimensional Poisson: Training and testing curves for different numbers of RFM basis functions.
Figure 54: *
Figure 55: *
Figure 56: *
Figure 57: High-Dimensional Poisson: Worst-case sample from the high-dimensional Poisson dataset. Visualization of $u(\mathbf{y})$ along the cross-section $(y_{3},y_{4})$ with all other coordinates fixed as $y_{i}=0$ for $i\neq 3,4$. (a) Pointwise comparison between the prediction and the ground truth. (b) Exact solution. (c) Predicted solution.
5 Conclusion
We propose a coefficient-to-coefficient (C2C) operator learning framework constructed on fixed basis spaces for both the input and output functions, termed the Fixed-Basis Coefficient-to-Coefficient Operator Network (FB-C2CNet). Unlike conventional neural operator frameworks that take pointwise samples on spatial grids as inputs, our method employs the coefficients associated with a set of pre-selected fixed basis functions. This representation substantially reduces the input dimensionality and the number of trainable parameters, thereby alleviating the training difficulty. Moreover, by fixing the basis functions a priori, the proposed framework improves both the stability and efficiency of the training process.
In this work, we employ two representative types of basis functions: the Random Feature Model (RFM) and the Finite Element Method (FEM). We investigate how different parameter settings influence the spectral properties of the coefficient space, showing that configurations with a higher effective rank and more balanced coefficient variations tend to facilitate training and enhance generalization. Our theoretical analysis further demonstrates that the accuracy of the proposed FB-C2CNet is ultimately bounded by the approximation capability of the chosen output function space $V_{\mathrm{out}}$.
We conduct extensive numerical experiments on a wide range of problems, including linear and nonlinear operators, regular and complex domains, and one-, two-, and high-dimensional settings. The results demonstrate the robustness and versatility of the proposed framework. Comparative studies show that the C2C formulation is easier to train and generalizes better than the traditional point-to-coefficient (P2C) approach. The proposed FB-C2CNet also exhibits strong resolution generalization, maintaining high accuracy across different discretization levels. Moreover, we extend the framework to multi-input/output operator learning, as demonstrated in the elastic plate and 2D Darcy flow problems, where multiple input and output functions are involved. The vector-valued RFM formulation in this setting achieves superior efficiency and accuracy, further validating the flexibility and effectiveness of the proposed approach.
Acknowledgments
The work was supported by the Project of Hetao Shenzhen-HKUST Innovation Cooperation Zone HZQB-KCZYB-2020083.
Appendix
Appendix A Experiments Details
Learning rate
Figure 58: Learning Rate Schedule. The learning rate decays from $10^{-3}$ to $10^{-6}$ for training stability, with a gradual decrease applied within every 10,000 epochs.
Parameter setting
In our method, the following parameters are primarily taken into consideration.
Table 2: Key quantities considered in this case.
Quantity Description
Number of training data $N_{\text{train}}$
Number of testing data $N_{\text{test}}$
Datasets Number of input grid points $n_{\text{in}}$
Number of output grid points $n_{\text{out}}$
Number of input basis functions $m_{\text{in}}$
Basis systems Number of output basis functions $m_{\text{out}}$
Initial Variance $R_{m}$
Random Feature Model Number of PoU $M_{p}$
Cut-off cutcut
Table 3: The parameters for each dataset and input (output) RFM basis systems.
$N_{\text{train}}$ $N_{\text{test}}$ $n_{in}$ $n_{out}$ $m_{in}$ $m_{out}$ $R_{m}$ $M_{p}$ cutcut
1D Darcy Flow 1000 800 2000 2000 128 32 3 (3) [8] ([4]) 1e-2
2D Darcy Flow (regular) 2400 600 19881 19881 1024 2048 0.1 (3) [8,8] ([8,8]) 1e-1
2D Poisson Equation 1600 400 1780 1780 256 400 3 (3) [2,2] ([8,8]) 1e-2
Elastic Plate Equation 1850 100 101 $1048\times 2$ 64 512 3 (3) [4] ([2,2]) 1e-3
2D Darcy Flow (Lshaped) $51000$ 7000 $736\times 2$ 450 1024 1024 3 (3) [16,16] ([8,8]) 1e-2
2D Darcy Flow (Lshaped) $5100$ 7000 $736\times 2$ 450 1024 1024 3 (3) [16,16] ([8,8]) 1e-2
2D Darcy Flow (Lshaped) $1000$ 7000 $736\times 2$ 450 1024 1024 3 (3) [16,16] ([8,8]) 1e-2
10D Poisson Equation 8 2 $31^{10}$ $31^{10}$ 128 128 0.3(0.3) [1,1,1,1,1,1,1,1,1,1] 1e-1
10D Poisson Equation 8 2 $31^{10}$ $31^{10}$ 8192 8192 0.3 (0.3) [1,2,1,1,2,1,2,2,1,2] 1e-1
The relative accuracy of our method is as followed.
Table 4: Mean Relative Error $L^{2}$ of the testing data.
FEM-C2C RFM-C2C
1D Darcy Flow 4.26e-2 1.20e-2
2D Darcy Flow (Regular) 7.04e-3 6.07e-3
2D Poisson Equation 8.41e-3 6.40e-3
Elastic Plate Equation 6.26e-1 8.92e-3
2D Darcy Flow (Lshaped) 2.94e-1 3.71e-2
10D Poisson Equation - 6.48e-3