### v2

# Learning Operators through Coefficient Mappings in Fixed Basis Spaces

Abstract
Operator learning has emerged as a promising paradigm for approximating solution operators of partial differential equations (PDEs). However, conventional approaches typically rely on pointwise function discretizations, which often suffer from the curse of dimensionality, mesh dependence, and prohibitive training costs in high-resolution settings. To address these challenges, we propose the Fixed-Basis Coefficient to Coefficient Operator Network (FB-C2CNet), a novel framework that learns the operator mapping within the coefficient spaces induced by prescribed, fixed basis functions. Unlike existing methods that learn basis functions dynamically or rely on extensive sensor grids, FB-C2CNet encodes input functions onto a fixed set of basis functions (such as random features or finite element bases) and employs a neural network to predict the expansion coefficients of the solution. By decoupling basis selection from network training, this formulation significantly reduces the dimensionality of the input-output spaces and the number of trainable parameters. We further introduce metrics such as effective rank to analyze how the spectral properties of the coefficient space influence generalization performance. Extensive numerical experiments across a wide spectrum of benchmarks—including linear, nonlinear, and high-dimensional problems—demonstrate that FB-C2CNet achieves competitive predictive accuracy while reducing training time by orders of magnitude compared to conventional neural operators.
keywords:
Operator Learning , Random Feature Method, Function Encoders , Basis Functions
journal: Computer Methods in Applied Mechanics and Engineering
[…]
11footnotetext: Author names are ordered alphabetically to indicate equal contribution.
1 Introduction
Deep learning has recently emerged as a powerful paradigm in scientific computing, with applications ranging from solving differential equations to accelerating large-scale simulations and data-driven modeling [37, 5, 35, 29]. Among various methodologies, operator learning has garnered significant attention, as it provides a principled framework for approximating mappings between infinite-dimensional function spaces directly. Unlike neural PDE solvers such as Physics-Informed Neural Networks (PINNs) [38, 27, 13], the Deep Ritz Method (DeepRitz) [16], Weak Adversarial Networks (WANs)  [45, 3], and random feature models (RFM) [10, 15, 41, 9], which approximate the solution of a partial differential equation (PDE) at finitely many spatial or temporal points, operator learning [33, 31, 34] aims to approximate the entire mapping from input functions such as initial conditions, boundary conditions, or source terms to output functions, typically the corresponding PDE solutions. For instance, consider the operator $G:(a,f)\mapsto u,$ which maps a coefficient function $a$ and a source term $f$ to the solution $u$ of the elliptic PDE: $\nabla\cdot(a\nabla u)=f,$ subject to appropriate boundary conditions. This example will be used as a running illustrative case throughout the paper. This operator-based perspective represents a paradigm shift: once trained, operator learning models can serve as fast and generalizable solvers applicable to a wide range of problem settings. This stands in contrast to conventional numerical algorithms, which typically rely on case-specific discretizations and require repeated, computationally intensive simulations for each new instance.
Despite its great potential, operator learning faces several fundamental challenges. One of the most critical issues lies in the discretization of infinite-dimensional function spaces. In other words, a key question is: what should serve as the input and output representations of a neural operator? To make the problem computationally tractable, these infinite-dimensional spaces are typically approximated by finite-dimensional subspaces. In most existing operator learning frameworks [33, 31, 34, 40, 44, 9], functions are discretized through their nodal values at a collection of collocation points, which are then used as the neural network input and output. While this pointwise encoding is straightforward and widely adopted, it suffers from several inherent drawbacks, including strong mesh dependence and degraded generalization performance as the spatial resolution increases. Moreover, in two- and three-dimensional problems, achieving adequate accuracy often requires a large number of grid points, leading to extremely high-dimensional inputs. This, in turn, results in neural networks with a prohibitively large number of parameters, significantly increased training cost, and long training times.
In this work, we propose a fixed basis coefficient-to-coefficient (FB-C2C) operator learning framework. The key idea is to represent both input and output functions in terms of chosen basis functions, and to train a neural network to learn the mapping between their expansion coefficients. Let $G:\mathcal{X}\rightarrow\mathcal{Y}$ be an operator that maps an input function $f\in\mathcal{X}$ to an output function $G(f)\in\mathcal{Y}$. We approximate the input function space $\mathcal{X}$ by a finite-dimensional subspace $V_{\mathrm{in}}=\text{span}\{\phi_{j}(x)\}_{j=1}^{m_{1}}$ and the output function space $\mathcal{Y}$ by $V_{\mathrm{out}}=\text{span}\{\psi_{j}(y)\}_{j=1}^{m_{2}}$. For each input function $f(x)$, we approximate it as
$f(x)\approx\sum_{j=1}^{m_{1}}a_{j}\phi_{j}(x),$
and use the coefficient vector $\bm{a}=(a_{1},a_{2},\dots,a_{m_{1}})$ as the input to a neural network $NN_{\theta}$, where $\theta$ denotes the trainable parameters. The network outputs the coefficient vector corresponding to the output basis functions, yielding an approximation of the operator output as
$G(f)(y)\approx\sum_{j=1}^{m_{2}}\big[NN_{\theta}(\bm{a})\big]_{j}\,\psi_{j}(y).$
In this way, the proposed framework learns the mapping directly between input and output coefficients, thereby realizing a fixed basis coefficient-to-coefficient (FB-C2C) operator learning paradigm that provides a compact representation and facilitates improved accuracy and generalization compared to nodal-value encodings. The main idea of our method is illustrated in Figure 1. A detailed description of the proposed methodology will be presented in Section 2.
Figure 1: Fixed-Basis Coeffiecient to Coeffiecient Operator Network (FB-C2CNet) framework.
This section proceeds with a review of the relevant literature, followed by a summary of the main contributions of this work.
1.1 Related work
Neural network–based methods for solving partial differential equations (PDEs) have been extensively studied in recent years. Existing approaches can be broadly categorized into two classes.
The first class focuses on direct approximation of PDE solutions, where neural networks are trained to represent individual solutions. Representative methods include PINNs [38, 27], the Deep Ritz method [16], WANs [45, 3], and RFM [10, 15]. These methods typically incorporate PDE constraints through loss functions derived from strong, weak, or variational formulations.
The second class is operator learning, which aims to approximate solution operators mapping input functions—such as coefficients, source terms, or boundary conditions—to solution functions. Representative approaches include DeepONet [33] and the Fourier Neural Operator (FNO) [31]. A variety of extensions have since been proposed to address multiscale problems, domain decomposition, and complex physical systems [19, 28, 30]. The present work is situated within the operator learning framework.
Operator Learning
DeepONet [33] is one of the earliest and most influential works on operator learning. Building on the Universal Approximation Theorem [22] for Operators, it introduces the framework:
$G(f)(y):=NN_{brach}(f)\cdot NN_{trunk}(y)=\underbrace{\sum_{k=1}^{p}\sum_{i=1}^{q}c_{i}^{k}\,\sigma\!\left(\sum_{j=1}^{m}\xi_{ij}^{k}\,f(x_{j})+\theta_{i}^{k}\right)}_{\text{Branch}}\;\underbrace{\sigma\!\left(\mathbf{w}_{k}\cdot\mathbf{y}+\zeta_{k}\right),}_{\text{Trunk}}$ (1)
where the Branch Net takes as input the function values at sensor points $f(x_{j})$, while the Trunk Net takes spatial coordinates and outputs basis functions. In this paper, we refer to such approaches as point-to-coefficient (P2C) methods, emphasizing that the Branch Net maps pointwise function evaluations to coefficient representations. The final output is obtained by combining the coefficients from the Branch Net with the basis from the Trunk Net. Following this, POD-DeepONet [34] replaced the Trunk Net with Proper Orthogonal Decomposition (POD) modes extracted from training data, while keeping the Branch Net to learn their coefficients. Later studies further explored alternative trunk representations, including fixed bases from Extreme Learning Machines (ELM) [40], finite element shape functions in Mesh-Informed Neural Networks (MINNs) [44], and random feature models with partition-of-unity strategies [9]. Despite these variations, the Branch Net still relies on $f(x_{j})$ as input. However, this design suffers from scalability issues: in higher dimensions, the number of sensor points becomes large, leading to high-dimensional inputs, larger networks, slower training, and potentially weaker generalization.
To address the challenges posed by high-dimensional inputs in operator learning, recent work has explored integrating neural networks with principled dimensionality reduction strategies [1, 4, 42, 20, 21, 46]. Among these, three approaches are particularly relevant to our work. The Basis Operator Network (BasisONet) [24] represents input functions using coefficients with respect to a prescribed basis. In this framework, the Branch Net takes the basis coefficients as input, while the Trunk Net encodes spatial locations, and both networks are trained jointly, with orthogonality constraints imposed on the Trunk Net throughout training. Another closely related approach is the Basis-to-Basis (B2B) operator learning method [25], in which the Trunk and Branch Nets are trained separately. In the Resolution-Independent Neural Operator framework [2], the trunk network selects basis functions from a predefined dictionary, and dictionary learning techniques are employed to adaptively construct suitable continuous basis functions.
A common limitation of these approaches is that the underlying basis functions are learned during training, which can be computationally expensive and potentially unstable. In contrast, our approach directly selects a predefined function space—such as finite element bases or random feature models—as the function encoder, and computes the corresponding expansion coefficients via a least-squares projection, which we refer to as the encoding step. These coefficients are then used as inputs to the neural operator, leading to improved training stability and reduced overall computational cost.
1.2 Contributions.
In this work, we propose the Fixed-Basis Coefficient Operator Network (FB-C2CNet), a new operator learning framework built upon fixed basis function systems. Our main contributions are summarized as follows:
  • 1.
    We introduce FB-C2CNet, which decouples operator learning into two stages: projection onto a fixed basis and learning in the coefficient space. This design eliminates the need to train the basis or encoder, thereby significantly reducing training cost.
  • 2.
    We investigate different choices of approximation function spaces, including random feature models and finite element bases, and study how the resulting coefficient representations influence the generalization performance of neural operators. In particular, we introduce effective rank and coefficient variation to quantify whether the distribution of the input coefficients is conducive to learning representations that enable robust generalization.
  • 3.
    We validate the effectiveness of the proposed FB-C2CNet framework through extensive numerical experiments across a diverse set of benchmark problems spanning multiple PDE classes, including steady-state elliptic equations, time-dependent nonlinear PDEs, Helmholtz equations with only weak solutions, and inverse boundary problems. Across all benchmarks, FB-C2CNet consistently achieves competitive or superior accuracy compared to existing neural operator methods, while requiring substantially reduced training cost.
  • 4.
    We further extend the proposed framework to multi-input/output operator learning, where mappings are defined between multiple input and output functions. We compare scalar-valued coefficient representations with vector-valued constructions based on shared vector-valued basis functions, and demonstrate that the RFM-based vector-valued formulation consistently achieves improved stability and accuracy. Moreover, an inverse boundary problem for the Stokes system is used as a representative application, illustrating the effectiveness of the proposed framework in realistic inverse problem settings.
The remainder of the paper is organized as follows. Section 2 provides a detailed introduction to the proposed fixed basis coefficient-to-coefficient operator learning framework. Section 3 discusses the design choices of key components, including the selection of basis function spaces and strategies for coefficient computation. Section 4 presents numerical experiments demonstrating the effectiveness, efficiency, and accuracy of the proposed method, followed by conclusions and future directions in Section 5.
2 Methodology
In this work, our goal is to approximate the mapping $G:\mathcal{X}\mapsto\mathcal{Y}$ using neural networks, where $\mathcal{X}$ and $\mathcal{Y}$ denote infinite-dimensional spaces of real-valued functions defined on a bounded domain in $\mathbb{R}^{d}$. The operator $G$ maps an input function $f\in\mathcal{X}$ to the corresponding solution $u\in\mathcal{Y}$ of a given partial differential equation (PDE). Our approach focuses on the representation of the input function $f$. Instead of adopting a pointwise discretization, we employ a basis-function representation: the input function $f$ is encoded by its expansion coefficients with respect to a chosen set of basis functions, and the neural network is trained to predict the coefficients associated with the basis representation of the solution. Compared with pointwise discretization, using coefficients as the input encoding significantly reduces the dimensionality of the neural network input. This reduction in dimension decreases the number of network parameters and leads to faster convergence during training, particularly in two- or higher-dimensional settings. In addition, the coefficient-based representation provides a more compact and structured description of the function space, mitigates the redundancy inherent in dense pointwise sampling, improves numerical stability, and enhances interpretability by directly linking the learned coefficients to the underlying basis functions. It is worth emphasizing that the basis functions in our framework are pre-selected and fixed, rather than learned during training. This design makes the method more efficient and results in faster and more stable training.
The remainder of this section provides a detailed introduction to our method and elaborates on its specific components.
2.1 Problem formulation.
Let $\mathcal{X},\mathcal{Y}$ be function spaces, our goal is to use neural network to learn the operator
$G:\mathcal{X}\to\mathcal{Y}.$ (2)
Suppose we are given the training data in the form of input-output pair $\{f^{(k)},u^{(k)}\}_{k=1}^{M}\subset\mathcal{X}\times\mathcal{Y}$, where $f^{(k)}\in\mathcal{X}$ is usually the source term or initial condition for the PDE problem, and $u^{(k)}\in\mathcal{Y}$ is the solution of the PDE. To numerically resolve the PDE, a set of collocation points must be specified within the domain. The input functions are evaluated at the points $(\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{n_{1}})$ and the output functions at $(\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{y}_{n_{2}})$, respectively. Thus, the training data set becomes: $\{\mathbf{f}^{(k)}=[f^{(k)}(\mathbf{x}_{1}),f^{(k)}(\mathbf{x}_{2}),\dots,f^{(k)}(\mathbf{x}_{n_{1}})]^{T},\mathbf{u}^{(k)}=[u^{(k)}(\mathbf{y}_{1}),u^{(k)}(\mathbf{y}_{2}),\dots,u^{(k)}(\mathbf{y}_{n_{2}})]^{T}\}_{k=1}^{M}$. In practice, we approximate this mapping using a neural network $NN_{\bm{\theta}}$, where $\bm{\theta}$ represents the neural network’s parameters. In the following, we introduce our proposed method, termed FB-C2CNet.
2.2 Fixed-Basis Coefficient to Coefficient Operator Network (FB-C2CNet)
Our method consists of two main stages. The first stage is fixed basis function selection. In this stage, we specify two basis function systems, denoted as $V_{\text{in}}=\text{span}\{\phi_{i}(\mathbf{x})\}_{i=1}^{m_{1}}$ and $V_{\text{out}}=\text{span}\{\psi_{i}(\mathbf{y})\}_{i=1}^{m_{2}}$, which are employed to approximate the input function space $\mathcal{X}$ and the output function space $\mathcal{Y}$, respectively. In the subsequent step, a given input function $f(\mathbf{x})\in\mathcal{X}$ is represented in terms of the basis $V_{\text{in}}=\text{span}\{\phi_{i}(\mathbf{x}),\phi_{2}(\mathbf{x}),\dots,\phi_{m_{1}}(\mathbf{x})\}$, i.e.,
$f(\mathbf{x})\approx\sum_{j=1}^{m_{1}}a_{j}\,\phi_{j}(\mathbf{x}),\quad\phi_{j}\in V_{\text{in}},f(\mathbf{x})\in\mathcal{X},$ (3)
where $\mathbf{a}=[a_{1},a_{2},\dots,a_{m_{1}}]\in\mathbb{R}^{m_{1}}$ are the expansion coefficients associated with $f$. The expansion coefficients $\mathbf{a}$ are obtained by solving the following regularized least-squares problem:
$\min_{\mathbf{a}\in\mathbb{R}^{m_{1}}}\;\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\left|f(\mathbf{x}_{i})-\sum_{j=1}^{m_{1}}a_{j}\,\phi_{j}(\mathbf{x}_{i})\right|^{2}+\lambda\|\mathbf{a}\|_{2}^{2},$ (4)
where $\{x_{i}\}_{i=1}^{n_{1}}$ are the sampling points in the input domain, and the second term provides $L^{2}$-regularization to stabilize the coefficient estimation. In the later sections of this paper, we will provide a more detailed discussion on the computation of the coefficients and analyze how different solution strategies may influence the performance of FB-C2CNet.
Following the above procedure, we obtain the coefficient representations $\{\mathbf{a}^{(k)}=[a^{(k)}_{1},a^{(k)}_{2},\dots,a^{(k)}_{m_{1}}]\}_{k=1}^{M}$, where each $\mathbf{a}^{(k)}\in\mathbb{R}^{m_{1}}$ corresponds to the expansion coefficients of the input function $f^{(k)}$ with respect to the input basis function systems $V_{\text{in}}$. We then take these coefficient representations as the input to our neural operator. This constitutes the second stage of our framework, in which a neural network is employed to approximate the underlying operator that maps the input coefficients $\mathbf{a}^{(k)}\in\mathbb{R}^{m_{1}}$ to the corresponding output coefficients $\mathbf{b}^{(k)}\in\mathbb{R}^{m_{2}}$ associated with the output basis function systems $V_{\text{out}}$. Specifically, the neural network $NN_{\bm{\theta}}$ is trained to approximate the mapping between the input and output coefficient representations, i.e.,
$NN_{\bm{\theta}}:\mathbf{a}\mapsto\mathbf{b},$ (5)
where $\mathbf{a}\in\mathbb{R}^{m_{1}}$ and $\mathbf{b}\in\mathbb{R}^{m_{2}}$ denote the coefficient representations of the input and output functions with respect to the fixed bases $V_{\text{in}}$ and $V_{\text{out}}$, respectively. In practice, we adopt a fully connected neural network (FNN) as the architecture for the operator approximation. Therefore, the overall structure of our proposed FB-C2CNet can be summarized as follows:
[…]
Here the set of trainable parameters in FB-C2CNet is denoted by $\bm{\theta}=\{c_{i}^{k},\,\xi_{ij}^{k},\,\eta_{i}^{k}\}$, which collectively define the weights and biases of the underlying fully connected network. $m_{1}$ denotes the number of basis functions in $V_{\text{in}}$, while $m_{2}$ denotes the number of basis functions in $V_{\text{out}}$. In the following, we present the training procedure adopted for FB-C2CNet.
Training method
The loss function employed to train FB-C2CNet is the mean relative $L^{2}$ error, defined as
$L(\bm{\theta})=\frac{1}{M}\sum_{k=1}^{M}\frac{\left\|NN_{\theta}(\mathbf{a}^{(k)})\cdot\mathbf{\Psi}-\mathbf{u}^{(k)}\right\|_{2}}{\left\|\mathbf{u}^{(k)}\right\|_{2}},$ (7)
where $\{\mathbf{a}^{(k)}=[a_{1}^{(k)},a_{2}^{(k)},\dots,a_{m_{1}}^{(k)}]\}_{k=1}^{M}$ denotes the set of input coefficient vectors, $\{\mathbf{u}^{(k)}=[u^{(k)}(\mathbf{y}_{1}),u^{(k)}(\mathbf{y}_{2}),\dots,u^{(k)}(\mathbf{y}_{n_{2}})]\}_{k=1}^{M}$ represents the corresponding sampled outputs, and $\mathbf{\Psi}\in\mathbb{R}^{n_{2}\times m_{2}}$ is the basis evaluation matrix with entries
$\Psi_{ij}=\psi_{j}(\mathbf{y}_{i}),\quad 1\leq i\leq n_{2},\;1\leq j\leq m_{2}.$ (8)
Here, $NN_{\bm{\theta}}$ is the neural operator parameterized by $\bm{\theta}$.
We adopt the mean relative error loss instead of the plain mean squared error (MSE) in order to normalize the approximation error with respect to the magnitude of the ground truth. This normalization ensures scale-invariance across different training samples: functions with larger amplitudes do not dominate the optimization process, while functions with smaller amplitudes are not overlooked. Such a formulation is particularly important in operator learning, where the outputs $\{\mathbf{u}^{(k)}\}_{k=1}^{M}$ may vary significantly in scale across different samples. By employing the relative loss, FB-C2CNet achieves a more balanced training process and yields a more robust generalization performance. The overall procedure is illustrated in Figure 2 and Algorithm. For clarity, the main notation used throughout the paper is summarized in Table 1.
Figure 2: Workflow of the proposed Fixed-Basis Coefficient to Coefficient Network (FB-C2CNet), exemplified via RFM Basis functions. It consists of three main stages: (i) fixed basis function selection, (ii) coefficient representation of input and output functions, and (iii) coefficient-to-coefficient operator learning via a neural network.
Table 1: Summary of notation and key quantities used in the paper.
Category Description Notation
Dataset Number of training samples $N_{\mathrm{train}}$
Number of testing samples $N_{\mathrm{test}}$
Number of input grid points $n_{1}$
Number of output grid points $n_{2}$
Basis representation Number of input basis functions $m_{1}$
Number of output basis functions $m_{2}$
Random Feature Model (RFM) Initial variance of random features $R_{m}$
Number of partitions of unity (PoU) $M_{p}$
Spectral cut-off parameter $\mathrm{cut}$
Algorithm Fixed Space Coefficient-to-Coefficient Operator Learning
Phase I: Choose fixed input/output basis systems: $V_{\mathrm{in}}=\text{span}\{\phi_{j}(\bm{x})\}_{j=1}^{m_{1}}$,  $V_{\mathrm{out}}=\text{span}\{\psi_{j}(\bm{y})\}_{j=1}^{m_{2}}$.
Strategy 1: Random feature model
Input: $m_{1},m_{2}$, $R_{m}^{\mathrm{in/out}}$, $M_{p}^{\mathrm{in/out}},J_{n}^{\mathrm{in/out}}$
Procedure: 1. For $n=1,\dots,M_{p}^{\mathrm{in}}$ and $j=1,\dots,J_{n}^{\mathrm{in}}$: $\phi_{nj}(\bm{x})=\omega_{n}(\bm{x})\sigma(\bm{k}^{\mathrm{in}}_{nj}\!\cdot\!\bm{x}+b^{\mathrm{in}}_{nj}),$ with $\bm{k}^{\mathrm{in}}_{nj}\!\sim\!U([-R_{m}^{\mathrm{in}},R_{m}^{\mathrm{in}}]^{d})$, $b^{\mathrm{in}}_{nj}\!\sim\!U([-R_{m}^{\mathrm{in}},R_{m}^{\mathrm{in}}])$. 2. Similarly construct $\psi_{nj}(\bm{y})$ for output basis with $\bm{k}^{\mathrm{out}}_{nj}$, $b^{\mathrm{out}}_{nj}$. Output: $V_{\mathrm{in}}=\text{span}\{\phi_{j}(\bm{x})\}_{j=1}^{m_{1}}$, $V_{\mathrm{out}}=\text{span}\{\psi_{j}(\bm{y})\}_{j=1}^{m_{2}}$.
Strategy 2: Finite element model
Input: element types $E^{\mathrm{in/out}}$, polynomial degrees $p^{\mathrm{in/out}}$, evaluation points $\bm{\xi}^{\mathrm{in/out}}$
Procedure: 1. Generate the mesh $\mathcal{T}_{h}$; then, for the chosen element family and degree, construct the global FE basis. Output: $V_{\mathrm{in}}=\text{span}\{\phi_{j}(\bm{x})\}_{j=1}^{m_{1}}$, $V_{\mathrm{out}}=\text{span}\{\psi_{j}(\bm{y})\}_{j=1}^{m_{2}}$.
Phase II: Encode input coefficients via least squares.
Input:data sets $\{\mathbf{f}^{(k)}\}_{k=1}^{M}$, each $\mathbf{f}^{(k)}=\{f^{(k)}(\bm{x}_{i})\}_{i=1}^{n_{1}}$; penalty $\lambda$
for $k=1$ to $M$ do
  compute $\bm{a}^{(k)}\in\mathbb{R}^{m_{1}}$ by
   $\bm{a}^{(k)}=\underset{\bm{a}\in\mathbb{R}^{m_{1}}}{\arg\min}\;\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\left|f^{(k)}(\bm{x}_{i})-\sum_{j=1}^{m_{1}}a_{j}\phi_{j}(\bm{x}_{i})\right|^{2}+\lambda\|\bm{a}\|_{2}^{2}.$
end for
[…]
Phase III: Train coefficient-to-coefficient network $NN_{\theta}$.
Input:targets $\{\mathbf{u}^{(k)}\}_{k=1}^{M}$, each $\mathbf{u}^{(k)}=\{u^{(k)}(\bm{y}_{i})\}_{i=1}^{n_{2}}$; step size $\eta$
form $\Psi\in\mathbb{R}^{n_{2}\times m_{2}}$ with $\Psi_{ij}=\psi_{j}(\bm{y}_{i})$
while not converged do
  for each $k$:
    $\bm{b}^{(k)}=NN_{\theta}(\bm{a}^{(k)})$, $\hat{\bm{u}_{\bm{\theta}}}^{(k)}=\Psi\bm{b}^{(k)}$, $\bm{u}^{(k)}=\big(u^{(k)}(\bm{y}_{i})\big)_{i=1}^{n_{2}}$
  loss:
   $L(\bm{\theta})=\frac{1}{M}\sum_{k=1}^{M}\frac{\|\hat{\bm{u}_{\bm{\theta}}}^{(k)}-\bm{u}^{(k)}\|_{2}}{\|\bm{u}^{(k)}\|_{2}}$
  update: $\bm{\theta}\leftarrow\bm{\theta}-\eta\nabla_{\bm{\theta}}L(\bm{\theta})$
end while
Output: $N\!N_{\bm{\theta}}$
3 Basis Systems and Coefficient Estimation
The goal of operator learning is to approximate an underlying mapping $G:\mathcal{X}\mapsto\mathcal{Y},$ where both $\mathcal{X}$ and $\mathcal{Y}$ are infinite-dimensional function spaces. Directly learning such a map is intractable, so it is necessary to embed $\mathcal{X}$ and $\mathcal{Y}$ into finite-dimensional subspaces. In our proposed FB-C2CNet, this embedding is realized by choosing an input function system $\{\phi_{j}(\bm{x})\}_{j=1}^{m_{1}}$ and an output function system $\{\psi_{j}(\bm{y})\}_{j=1}^{m_{2}}$, which span the subspaces
$V_{\mathrm{in}}=\mathrm{span}\{\phi_{j}\}_{j=1}^{m_{1}}\quad\text{and}\quad V_{\mathrm{out}}=\mathrm{span}\{\psi_{j}\}_{j=1}^{m_{2}}.$
Each source function $f\in\mathcal{X}$ and target function $u=G(f)\in\mathcal{Y}$ can then be projected onto these subspaces as
$f(\bm{x})\approx\sum_{j=1}^{m_{1}}a_{j}\phi_{j}(\bm{x}),\qquad u(\bm{y})\approx\sum_{j=1}^{m_{2}}b_{j}\psi_{j}(\bm{y}),$
where $\mathbf{a}=[a_{1},\dots,a_{m_{1}}]\in\mathbb{R}^{m_{1}}$ and $\mathbf{b}=[b_{1},\dots,b_{m_{2}}]\in\mathbb{R}^{m_{2}}$ are the corresponding coefficient representations. Hence, the infinite-dimensional operator learning problem is transformed into learning a finite-dimensional coefficient mapping
$NN_{\bm{\theta}}:\mathbb{R}^{m_{1}}\mapsto\mathbb{R}^{m_{2}},$
which approximates the coefficient-to-coefficient relationship induced by $G$. This is the core idea of our proposed Fixed-Basis Coefficient-to-Coefficient Operator Network (FB-C2CNet).
Therefore, in our proposed FB-C2CNet, two key aspects require particular attention: (i) the selection of the fixed basis function spaces, and (ii) the computation of the corresponding coefficients. In this section, we provide a detailed discussion of these two components.
3.1 Fixed Basis Systems
One of the key advantages of our proposed FB-C2CNet is that the input space $\mathcal{X}$ and the output space $\mathcal{Y}$ are directly represented using pre-specified basis functions. Since these basis functions (or the associated encoder) are fixed in advance, the network does not need to learn them during training. This design substantially lowers the training cost, as the learning process is reduced to estimating the coefficients with respect to the fixed bases, rather than simultaneously optimizing both the representation and the mapping. In this work, we primarily employ the random feature model and the finite element method as the basis function systems for constructing $V_{\text{in}}$ and $V_{\text{out}}$. In the following, we provide a brief introduction to these two basis function systems.
Random feature model
The random feature model (RFM) can be expressed as
$\phi(\bm{x};\bm{\theta})=\sum_{j=1}^{M}a_{j}\sigma(\bm{k}_{j}\cdot\bm{x}+b_{j}),$ (9)
where the inner parameters $\{\bm{k}_{j},b_{j}\}_{j=1}^{M}$ are randomly initialized, typically from a uniform distribution with $\bm{k}_{j}\sim U\left([-R_{m},R_{m}]^{d}\right)$ and $b_{j}\sim U\left([-R_{m},R_{m}]\right)$ [10]. The activation function $\sigma$ is often chosen as $\tanh(x)$ or $\sin(x)$. Importantly, only the outer coefficients $\{a_{j}\}_{j=1}^{M}$ are trainable, enabling the resulting optimization problem to be solved efficiently via standard linear algebra techniques such as least squares or QR decomposition. To further improve approximation, the partition of unity (PoU) method is often employed [10], leading to
$u(\bm{x})=\sum_{n=1}^{M_{p}}\omega_{n}(\bm{x})\sum_{j=1}^{J_{n}}a_{nj}\,\sigma(\bm{k}_{nj}\cdot\bm{x}+b_{nj}),$ (10)
where $\psi_{n}(\bm{x})$ are construction functions forming a PoU, and $\phi_{nj}(\bm{x})=\sigma(\bm{k}_{nj}\cdot\bm{x}+b_{nj})$ are random features. Here, $M_{p}$ denotes the number of partitions and $J_{n}$ the number of features in each partition.
In one dimension, $\omega_{n}$ can be either the characteristic function
$\omega_{n}^{a}(x)=\chi_{\{-1\leq\tilde{x}<1\}},$ (11)
or the smooth function
$\omega_{n}^{b}(x)=\begin{cases}\frac{1+\sin(2\pi\tilde{x})}{2},&-\tfrac{5}{4}\leq\tilde{x}<-\tfrac{3}{4},\\[6.0pt] 1,&-\tfrac{3}{4}\leq\tilde{x}<\tfrac{3}{4},\\[6.0pt] \frac{1-\sin(2\pi\tilde{x})}{2},&\tfrac{3}{4}\leq\tilde{x}<\tfrac{5}{4},\\[6.0pt] 0,&\text{otherwise},\end{cases}$ (12)
where $\tilde{x}=(x-x_{n})/r_{n}$ rescales the interval to [1,1][-1,1]. For $d$ dimensions, we set $\omega_{n}(\bm{x})=\prod_{k=1}^{d}\omega_{n}(x_{k}).$
A one-dimensional illustration is provided in Figure 3(a). This PoU-based RFM provides a flexible and efficient approximation framework for PDE solutions. Owing to its computational convenience, the random feature model has found applications across a broad range of problems [36, 11, 26]. Moreover, since it is a meshfree method, it can naturally handle complex computational domains without the need for mesh generation. In practice, it also exhibits advantages in high-dimensional problems, where its structure facilitates more tractable implementation and efficient approximation [18, 23, 32].
Finite element method
The finite element method (FEM) [7, 12] provides another natural choice of basis functions, particularly well-suited for approximating PDE solutions. Given a mesh $\mathcal{T}_{h}$ over the computational domain $\Omega\subset\mathbb{R}^{d}$, the FEM basis functions are constructed locally on each element. For example, in the case of linear ($P_{1}$) elements, the basis functions are the so-called hat functions, defined by
$\phi_{j}(\bm{x}_{i})=\delta_{ij},\quad 1\leq i,j\leq N_{h},$ (13)
where $\{\bm{x}_{i}\}_{i=1}^{N_{h}}$ are the nodal points of the mesh and $N_{h}$ is the number of degrees of freedom. The global FEM approximation of a function $u(\bm{x})$ can then be expressed as
$u_{h}(\bm{x})=\sum_{j=1}^{N_{h}}a_{j}\,\phi_{j}(\bm{x}),$ (14)
where the coefficients $\{a_{j}\}_{j=1}^{N_{h}}$ correspond to the nodal values of $u$. Higher-order elements ($P_{k}$, $k\geq 2$) can also be employed to improve approximation accuracy, in which case the basis functions are polynomials of degree $k$ defined on each element of the mesh. A one-dimensional illustration is provided in Figure 3(b). Due to their locality, conformity, and well-established approximation properties, FEM basis functions are particularly effective for incorporating boundary conditions and ensuring numerical stability. They have therefore become a standard tool for representing PDE solution spaces in scientific computing. Nevertheless, in high-dimensional settings they suffer from the curse of dimensionality, as the number of degrees of freedom grows exponentially with the dimension.
Other basis functions
In addition to FEM and random feature models, a variety of alternative basis systems can be used to construct approximation spaces. Classical spectral methods employ global orthogonal polynomials, such as Chebyshev or Legendre bases, and achieve exponential convergence for smooth solutions [6, 43]. Radial basis functions (RBFs), including Gaussian and multiquadric kernels, provide meshfree approximations well suited for scattered data and irregular domains [17, 8]. Wavelet bases offer multiresolution representations with simultaneous localization in space and frequency [14]. Each of these basis choices entails different trade-offs in accuracy, computational cost, and geometric flexibility, and can be incorporated into operator-learning frameworks as dictated by the problem setting.
(a) RFM basis.
(b) FEM basis.
Figure 3: Basis functions: (a) One-dimensional RFM example: $M_{p}=4$, with each partition containing $J_{n}=2$ random feature models, and random feature parameters sampled from the range $R_{m}=3$. (b) One-dimensional FEM example with $N_{h}=8$ nodal basis functions.
The choice of basis functions is problem-dependent and plays a crucial role in the accuracy and efficiency of the proposed framework. For simple one- and two-dimensional problems, finite element basis functions are often a natural and reliable choice, as they provide controllable accuracy and a solid theoretical foundation. For problems with complex geometries or irregular boundaries, finite element methods remain applicable, but constructing an appropriate mesh becomes highly non-trivial in practice, both in terms of the manual effort required and the associated computational cost. In such cases, mesh-free approaches based on random feature models offer a more flexible alternative. Moreover, for high-dimensional problems, the finite element method suffers from the curse of dimensionality, as the number of required basis functions grows exponentially with the spatial dimension. The random feature model, by contrast, avoids this issue and thus provides a more scalable and practical choice for high-dimensional operator learning. In the following numerical experiments, we present a series of cases comparing RFM-C2C and FEM-C2C, from which useful insights can be drawn regarding the choice of basis functions.
3.2 Coefficient representations and learnability
Once a fixed function space has been chosen, the next step is to determine the expansion coefficients $\{a_{j}\}_{j=1}^{m_{1}}$. At this stage, two fundamental questions naturally arise:
  • 1.
    What characterizes a desirable coefficient representation?
  • 2.
    How can the coefficients be computed efficiently and stably?
In the following, we address these questions in turn.
Coefficient Quality
We begin by examining the properties of coefficient representations that facilitate effective learning in the neural operator. The choice of coefficients is not merely a technical detail; it directly influences the trainability, convergence speed, and generalization ability of FB-C2CNet. For instance, poorly scaled or highly correlated coefficients may lead to ill-conditioned optimization landscapes, thereby making training unstable or inefficient. In contrast, well-structured coefficients that capture the essential variability of the target functions can significantly ease the learning burden of the network. In the following, we analyze what makes a coefficient representation “easy to learn,” and provide guidelines for constructing coefficient systems that are both expressive and numerically stable.
Suppose a set of basis functions $\{\phi_{j}(x)\}_{j=1}^{m_{1}}$ is chosen to span the input approximation space $V_{\mathrm{in}}$. Each training sample $f$ is projected onto this basis, yielding
$f(\mathbf{x})\approx\sum_{j=1}^{m_{1}}a_{j}\phi_{j}(\mathbf{x}),$
where $\bm{a}_{\mathrm{in}}=[a_{1},a_{2},\dots,a_{m_{1}}]^{\top}$ denotes the coefficient vector. Given a training dataset consisting of $M$ input functions $\{f^{(k)}(\mathbf{x})\}_{k=1}^{M}$, we obtain their corresponding coefficient representations $\{\bm{a}^{(k)}\}_{k=1}^{M}$, where
$\bm{a}^{(k)}=[a_{1}^{(k)},a_{2}^{(k)},\dots,a_{m_{1}}^{(k)}]^{\top}\in\mathbb{R}^{m_{1}\times 1}.$
Collecting all coefficient vectors into a matrix form gives
$A=\begin{bmatrix}a_{1}^{(1)}&a_{2}^{(1)}&\cdots&a_{m_{1}}^{(1)}\\[3.0pt] a_{1}^{(2)}&a_{2}^{(2)}&\cdots&a_{m_{1}}^{(2)}\\[3.0pt] \vdots&\vdots&\ddots&\vdots\\[3.0pt] a_{1}^{(M)}&a_{2}^{(M)}&\cdots&a_{m_{1}}^{(M)}\end{bmatrix}\in\mathbb{R}^{M\times m_{1}},$
Each row of $A$ corresponds to a projected training function in the coefficient space of $V_{\mathrm{in}}$. A useful diagnostic for assessing the learnability of such coefficient representations is the effective rank of the matrix $A$.
Definition 1 (Effective rank [39]).
For a matrix $\bm{A}\in\mathbb{R}^{m\times n}$ $(m>n)$ with singular values $\sigma_{1}\geq\sigma_{2}\geq\ldots\geq\sigma_{n}$, the truncation entropy of $\bm{A}$ is defined as:
$\operatorname{erank}(\bm{A})=\exp\left\{-\sum_{k=1}^{n}p_{k}\log p_{k},\right\},$ (15)
where $p_{k}=\frac{\sigma_{k}}{\|\bm{\sigma}\|_{1}}$, for $k=1,2,\ldots,n$, with $\bm{\sigma}=[\sigma_{1},\sigma_{2},\dots,\sigma_{n}]^{\top}$.
The effective rank can measure the intrinsic dimensionality of the coefficient distribution. Intuitively, a larger effective rank indicates that the coefficients span a richer and more balanced subspace of $\mathbb{R}^{m_{1}}$, which eases optimization and improves generalization. In contrast, when the effective rank is small, the coefficients concentrate in a low-dimensional subspace, leading to poor conditioning, slow convergence, and increased risk of overfitting.
In addition to the effective rank, the distribution of variations in the coefficient vector $\mathbf{a}$ associated with the basis functions $\{\phi_{j}(\mathbf{x})\}_{j=1}^{m_{1}}$ also plays an important role [47]. When the coefficients exhibit large fluctuations along only a few dominant basis directions, the learned model tends to overfit these components and generalizes poorly to unseen inputs. In contrast, when the variations in $\mathbf{a}$ are more evenly distributed among the basis functions, the representation becomes smoother and more balanced, leading to improved generalization performance.
Consequently, in the process of computing coefficients, our objective is to obtain a coefficient matrix that possesses both high effective rank and balanced variance distribution. Such properties not only enhance the expressiveness of the representation but also improve the trainability and generalization ability of FB-C2CNet.
Coefficient Computation
After fixing the basis functions, we proceed to the computation of the associated coefficients. Since the basis functions $\{\phi_{j}(\mathbf{x})\}_{j=1}^{m_{1}}$ are fixed and known, the Equation 4 can be solved directly in matrix form. Define the design matrix $\Phi\in\mathbb{R}^{n_{1}\times m_{1}}$ as
$\Phi_{ij}=\phi_{j}(\mathbf{x}_{i}),\quad 1\leq i\leq n_{1},\;1\leq j\leq m_{1},$ (16)
and let the input function values be $\mathbf{f}=\big[f(x_{1}),f(x_{2}),\dots,f(x_{n_{1}})\big]^{\top}\in\mathbb{R}^{n_{1}\times 1}.$ Then the coefficient vector $\mathbf{a}\in\mathbb{R}^{m_{1}\times 1}$ is determined from
$\Phi\mathbf{a}=\mathbf{f}.$ (17)
In our framework, the goal is to reduce the input dimensionality, and thus we typically encounter the case $m_{1}\ll n_{1}$, where the number of basis functions is much smaller than the number of sampling points. In this case, we compute $\mathbf{a}$ by solving the regularized least-squares problem
$\mathbf{a}^{\ast}=\arg\min_{\mathbf{a}\in\mathbb{R}^{m_{1}}}\|\Phi\mathbf{a}-\mathbf{f}\|_{2}^{2}+\lambda\|\mathbf{a}\|_{2}^{2},$ (18)
which admits the closed-form solution
$\mathbf{a}^{\ast}=(\Phi^{\top}\Phi+\lambda I)^{-1}\Phi^{\top}\mathbf{f}.$ (19)
The regularization parameter $\lambda$ controls the smoothness of the recovered coefficients: a larger $\lambda$ enforces stronger regularity and suppresses small-scale variations in $\mathbf{a}$.
Alternatively, the same regularization effect can be realized through the singular value decomposition (SVD). Let $\Phi=U\Sigma V^{\top}$ be the SVD of $\Phi$, with singular values $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{m_{1}}$. A truncated SVD (or cut-off) approach constructs the pseudo-inverse by discarding singular values below a threshold $\sigma_{\text{cut}}$, yielding
$\mathbf{a}^{\ast}=V\Sigma^{\dagger}U^{\top}\mathbf{f},$ (20)
where $\Sigma^{\dagger}=\mathrm{diag}(\sigma_{1}^{-1},\dots,\sigma_{r}^{-1},0,\dots,0)$ and $r$ is the number of retained singular modes. The cut-off parameter $\sigma_{\text{cut}}$ plays a role analogous to the regularization parameter $\lambda$: a larger cut-off (corresponding to stronger regularization) removes more high-frequency components, leading to smoother but less detailed coefficient representations.
From a spectral perspective, increasing the cut-off threshold effectively reduces the contribution of small singular values, resulting in a higher effective rank of the matrix $\Phi$ and reduced variations in the coefficient vector $\mathbf{a}$. Hence, the choice of $\sigma_{\text{cut}}$ (or equivalently, $\lambda$) directly controls the balance between stability, smoothness, and the expressive richness of the learned representation.
We observe that when the random feature model (RFM) is used as the basis, the cut-off strategy has a particularly significant impact on the resulting coefficients. This sensitivity arises because random features often lead to highly correlated basis functions, causing the singular value spectrum of $\Phi$ to decay rapidly [9]. As a result, the choice of truncation threshold strongly influences the quality of the coefficient representation.
Example
1D Darcy flow as an example to illustrate our finding.
To gain further intuition, we conduct a simple experiment to illustrate the effect of the SVD cut-off on the effective rank of the coefficient matrix $A$ in the RFM setting. Specifically, we consider the one-dimensional Darcy flow problem introduced in [27] (Details can be found in Section 4.1). For this test, we set the number of random features in both the input and output basis systems to $m_{1}=m_{2}=128$, and discretize the domain with $1000$ uniformly spaced grid points. Figure 4 shows the resulting coefficient distributions under different cut-off thresholds. We observe that as the cut-off increases, the coefficients become more concentrated, the variance across different modes decreases, and the effective rank correspondingly grows. This confirms that in the RFM setting, a larger cut-off can paradoxically lead to coefficient matrices with higher effective rank, owing to the rapid decay of the singular value spectrum.
Figure 4: Coefficient distributions. Illustration of the geometric distributions of the coefficient matrix obtained under different SVD cut-off thresholds in the RFM setting. For each cut-off value, we plot the histogram of nearest-neighbor distances in the high-dimensional coefficient space, where each bar corresponds to the empirical distribution of Euclidean distances between one sample and its closest distinct neighbor among the $3200$ samples in $A\in\mathbb{R}^{3200\times 300}$. As the cut-off threshold increases, the coefficients become more concentrated, the variance across modes decreases, and the effective rank correspondingly grows.
As shown in Figure 5, although the training curves are nearly identical across different cut-off thresholds, the generalization performance exhibits a clear dependence on the effective rank of the coefficient matrix. When the cut-off is small, the resulting coefficients yield a low effective rank, indicating that the representation is concentrated in a limited number of directions. Such low-rank structure restricts expressiveness and leads to poor generalization. By contrast, increasing the cut-off raises the effective rank, producing coefficient representations that span a richer subspace of the function space. This increase in effective rank is consistently associated with improved generalization performance, demonstrating that effective rank is a key indicator of generalization in FB-C2CNet. However, overly large cut-off values are also undesirable, as they overly concentrate coefficients and discard important structural information in the data.
Figure 5: Train and test curve. Illustration of the training and testing performance of FB-C2CNet when using coefficients computed with different SVD cut-off thresholds as inputs. The results show that the choice of cut-off significantly influences the coefficient representation, and consequently the training dynamics and generalization ability of the network.
3.3 Output-Space-Induced Error Analysis
In the following, we demonstrate that the ultimate accuracy of the proposed method is fundamentally constrained by the approximation capability of the output space $V_{\mathrm{out}}$. Let $u(\bm{y})$ denote the ground-truth solution of the target operator $G:\mathcal{X}\to\mathcal{Y}$, i.e., $u(\bm{y})\in\mathcal{Y}$. Once the output function system is fixed as $V_{\mathrm{out}}=\mathrm{span}\{\psi_{j}(\bm{y})\}_{j=1}^{m_{2}},$ any $u(\bm{y})\in\mathcal{Y}$ admits an approximation
$\hat{u}(\bm{y})=\sum_{j=1}^{m_{2}}b_{j}\psi_{j}(\bm{y}),\qquad\hat{u}\in V_{\mathrm{out}},$ (21)
where $\bm{b}=[b_{1},b_{2},\dots,b_{m_{2}}]$ denotes the output coefficient vector. The FB-C2CNet is trained to learn the coefficient-to-coefficient mapping
$NN_{\theta}:\bm{a}\in\mathbb{R}^{m_{1}}\mapsto\bm{b}\in\mathbb{R}^{m_{2}},$
so that the reconstructed prediction reads
$\hat{u}_{\bm{\theta}}(\bm{y})=NN_{\bm{\theta}}(\bm{a})\cdot\Psi,\qquad\Psi=[\psi_{1}(\bm{y}),\psi_{2}(\bm{y}),\dots,\psi_{m_{2}}(\bm{y})]\in\mathbb{R}^{n_{2}\times m_{2}}.$ (22)
We then consider the following triangle inequality:
$\|u-\hat{u}\|_{L^{2}}\;\leq\;\underbrace{\|\hat{u}-\hat{u}_{\theta}\|_{L^{2}}}_{\text{learnable operator error}}+\underbrace{\|u-\hat{u}_{\theta}\|_{L^{2}}}_{\text{error for the FB-C2CNet}}.$ (23)
Here, the first term $\|\hat{u}-\hat{u}_{\theta}\|_{L^{2}}$ measures the learnable discrepancy introduced by the neural mapping $NN_{\theta}$, indicating how well the network reproduces the coefficient-to-coefficient relation. The second term $\|u-\hat{u}_{\theta}\|_{L^{2}}$ represents the overall approximation error of the FB-C2CNet with respect to the reference solution $u$, which includes both the learnable error and the representation error associated with $V_{\mathrm{out}}$.
We take the lower bound of the left-hand side and denote it as the intrinsic limitation imposed by the output space $V_{\mathrm{out}}$, which is characterized by the projection error:
$\varepsilon_{\mathrm{proj}}:=\inf_{\mathbf{b}\in\mathbb{R}^{m_{2}}}\bigg\|u(\bm{y})-\sum_{j=1}^{m_{2}}b_{j}\psi_{j}(\bm{y})\bigg\|_{L^{2}},$ (24)
which satisfies $\varepsilon_{\mathrm{proj}}\leq\|u-\hat{u}\|_{L^{2}}$ for any choice of coefficients. This quantity can be regarded as a measure of the approximation capability of $V_{\mathrm{out}}$ to represent functions in $\mathcal{Y}$. In practice, the coefficients $\bm{b}$ are obtained via a least-squares projection of the target function onto the subspace $V_{\mathrm{out}}$. Combining with the above triangle inequality yields
$\varepsilon_{\mathrm{proj}}\;\leq\;\|\hat{u}-\hat{u}_{\theta}\|_{L^{2}}+\|u-\hat{u}_{\theta}\|_{L^{2}}.$ (25)
Therefore, under the assumption that the FB-C2CNet is sufficiently well trained such that the learnable mapping $NN_{\theta}$ accurately captures the underlying operator $G$, the total approximation error is dominated by and bounded by the projection error associated with the output space $V_{\mathrm{out}}$, i.e.,
$\varepsilon_{\mathrm{proj}}\;\leq\;\|u-\hat{u}_{\theta}\|_{L^{2}}.$ (26)
In the following numerical experiments, we will further verify that this theoretical prediction is consistent with the empirical results.
4 Numerical Experiments
In this section, we present a series of experiments across a broad spectrum of PDE settings to demonstrate the efficiency and effectiveness of our method: (i) the Darcy flow problem on regular domains in both one and two dimensions, (ii) the Poisson equation on a two-dimensional complex domain, (iii) 2D Darcy flow on a complex domain, (iv) elastic plate problems, (v) time dependent nonlinear KdV–Burgers equation, (vi) Helmholtz equation with only one weak solution, (vii) high-dimensional Poisson equations on regular domains. Inverse boundary problem for the Stokes system will be presented in next section. These examples encompass diverse dimensionalities, domain geometries, and operator learning scenarios. Datasets for experiments (i-v) follow [24, 25], data setups for other examples are described separately in their respective subsections.
It is worth noting that both the elastic plate and 2D Darcy flow problems in complex domains belong to the category of multi-input/output operator learning. In the elastic plate case, a single input function (the load distribution) is mapped to two output functions representing displacement components, while in the 2D Darcy flow case, two input functions (the permeability field and source term) are mapped to a single output pressure field. In these settings, we generalize the proposed framework to handle multi-component mappings and compare two encoding strategies: a scalar-valued representation formed by concatenating the coefficients of multiple inputs, and a vector-valued representation that employs shared vector-valued basis functions to jointly encode all components.
Unless otherwise stated, the neural operator is implemented as a fully connected network with architecture $m_{1},-,m_{2}]$, where $m_{1}$ and $m_{2}$ are the input and output dimensions determined by the number of basis functions used to approximate the input function space and output function space, and – represents a cascade of hidden layers. The depth and width of these hidden layers are not fixed; instead, they are designed to maximize representational capacity under a parameter budget of 500,000. All models are trained with Adam optimizer, employing a learning rate schedule that decays from $10^{-2}$ to $10^{-4}$ (as illustrated in Figure 25). All experiments were performed on a single NVIDIA RTX 4090 GPU. We utilize the Gmsh Python library for mesh generation. The resulting mesh is then used to construct the FEM linear basis functions. Additional implementation details and hyperparameter settings are provided in the Appendix.
4.1 1D Darcy Flow
In this example, we aim to learn the nonlinear Darcy operator for a one-dimensional system. A variant of the nonlinear 1D Darcy equation is given by
$\frac{d}{dx}\left(a(u)\frac{du}{dx}\right)$ $=f(x),\quad x\in(0,1),$ (27)
$u(0)=u(1)$ $=0,$
where the solution-dependent permeability is $a(u(x))=0.2+u^{2}(x)$ and the input term $f$ is modeled as a Gaussian random field, $f\sim\mathcal{GP}(0,k(x,x^{\prime}))$, with covariance kernel $k(x,x^{\prime})=\sigma^{2}\exp\left(-\frac{|x-x^{\prime}|^{2}}{\ell^{2}}\right)$, $\ell=0.04$, and $\sigma=1.0$. The dataset is from [25, 2]. The training set consists of 800 samples and the testing set consists of 200 samples. The spatial domain $[0,1]$ is discretized uniformly with 2000 grid points. Our objective is to learn the map:
$G:f(x)\mapsto u(x).$
C2C vs. P2C: Efficiency and Accuracy. In this example, we compare the coefficient-to-coefficient (C2C) and point-to-coefficient (P2C) approaches using basis functions from the random feature model and the finite element method. For the P2C setting, the network architecture and training procedure remain identical to those of the C2C framework, except that the neural network takes pointwise function values as inputs instead of coefficients. It means that the input layer dimension of neural network in the P2C method is equal to the number of spatial grid points ($n_{1}=n_{2}=2000$ in this case), while the C2C method uses a significantly smaller input dimension determined by the number of basis functions ($m_{1}=m_{2}=128$ in this case). This design highlights the effectiveness of using coefficient representations as neural network inputs. As shown in Figure 6, with 2000 spatial points, the P2C method requires a high-dimensional input (dimension 2000), resulting in a larger number of network parameters and more difficult training, which leads to slower loss decay and lower accuracy. In contrast, the proposed C2C method uses far fewer parameters and achieves higher accuracy, highlighting its efficiency.
(a)RFM basis.
(b) FEM basis.
Figure 6: 1D Darcy Flow: Training curve comparison between the C2C and P2C methods with different basis functions. The number of grid points is set to $n_{1}=n_{2}=2000$, and the number of basis functions is $m_{1}=128$ and $m_{2}=128$.
Generalization through different resolutions. From the original set of $n_{1}=n_{2}=2000$ points, we take every 20th point, resulting in a subset of 100 points used for training. We train both RFM-C2C and FEM-C2C on data with $n_{1}=n_{2}=100$ grid points, and then evaluate their performance on test datasets with $n_{1}=n_{2}=40$, $500$, $1000$, and $2000$ grid points, respectively. This experiment examines the ability of the learned operators to generalize across different spatial resolutions. The results, summarized in Figure 7, show that the proposed methods maintain stable accuracy across unseen resolutions and RFM-C2C achieves better performance.
Figure 7: 1D Darcy Flow: FB-C2CNet is trained on a uniform $100$-point grid and tested on different resolutions ($n_{1}=n_{2}=40,500,1000,2000$). The network exhibits strong resolution invariance with consistent test accuracy across all cases.
Table 2: 1D Darcy Flow: The comparison of the C2C framework and the P2C framework in two different RFM and FEM basis spaces. The cost time for the encoder means the time required to generate the basis functions and project the input functions to the corresponding basis space. The training time refers to the time required to train the neural operator.
Cost time Training time Test relative $L^{2}$ error Neural network structure
Basis Encoder C2C P2C C2C P2C C2C P2C
[…]
4.2 Darcy Flow in a rectangular domain
In this example, we consider a two-dimensional Darcy flow equation in the square domain $\Omega=(0,1)^{2}$. It is defined as
$-\nabla\cdot\left(a(x,y)\nabla u(x,y)\right)$ $=f(x,y),\quad(x,y)\in\Omega,$ (28)
$u(x,y)$ $=0,\quad(x,y)\in\partial\Omega,$
where $a(x,y)$ is the diffusion coefficient and $f(x,y)$ is the forcing function. Here we set $f(x,y)\equiv 1$. Our objective is to learn the map:
$G:a(x,y)\mapsto u(x,y).$
The diffusion coefficient $a(x,y)$ is sampled from random field $exp(\psi)$, where $\psi=\mathcal{N}(0,(-\Delta+9I)^{-2})$. The dataset is from [31, 24]. The training set consists of 2400 samples, and the testing set consists of 600 samples. The spatial domain $0,1]\times[0,1]$ is discretized uniformly with $141\times 141$ grid points.
C2C vs. P2C: Efficiency and Accuracy. In this example, similar to the 1D Darcy flow case, we compare the C2C and P2C approaches with different choices of basis functions. The learning rate for them starts from 1e31e-3 and is multiplied by 0.9 every 400 steps for a total of 10000 training steps. As shown in Figure 8 and Table 3, the results indicate that the C2C framework achieves lower training time since it has fewer parameters to be trained than the P2C framework, which in turn leads to higher accuracy in both basis functions. For the encoder part, generating FEM basis functions on a 2D grid is more computationally expensive than in 1D, making the efficiency advantage of the RFM basis even more pronounced in this case.
(a) RFM basis.
(b) FEM basis.
Figure 8: Darcy Flow in a rectangular domain: Training curve comparison between the C2C and the P2C frameworks with different basis functions. The number of grid points is set to $n_{1}=n_{2}=141\times 141$. The numbers of basis functions are $(m_{1},m_{2})=(512,2048)$ for the RFM basis, and $(m_{1},m_{2})=(553,2059)$ for the FEM basis.
Table 3: Darcy Flow in a rectangular domain: The comparison of the C2C framework and the P2C framework in two different RFM and FEM basis spaces. The cost time for the encoder means the time required to generate the basis functions and project the input functions to the corresponding basis space. The training time refers to the time required to train the neural operator.
Cost time Training time Test relative $L^{2}$ error Neural network structure
Basis Encoder C2C P2C C2C P2C C2C P2C
RFM 3.67s 274.40s 302.11s 6.74e-3 39.90e-3 [512,180,180,2048] $141^{2}$,180,180,2048]
FEM 447.05s 291.17s 317.65s 6.92e-3 36.00e-3 [553,180,180,2059] $141^{2}$,180,180,2059]
Comparison with different operator learning methods. For a comprehensive evaluation, we benchmark our approach against four representative operator‑learning methods: BasisONet, B2B, DeepONet, and FNO2D. Their training times (on the same training set), network parameter counts, and mean relative $L^{2}$ errors ($RL^{2}E$) / mean squared errors ($MSE$) (on the same test set) are reported in Table 4. In comparison with other methods, the two FB-C2C methods train significantly faster, except against DeepONet. However, DeepONet exhibits substantially larger errors. Although FNO2D achieves slightly lower errors than our method, it requires far more parameters and longer training times. Additional comparative results on other benchmark problems are provided in the Appendix.
Table 4: Darcy Flow in a rectangular domain: Performance comparison for 2D Darcy flow in a rectangular domain. For C2C method, the total time is the sum of the encoder time and the training time.
Time Params $RL^{2}E$ MSEMSE
BasisONet [24] 88m1s 2,712,174 7.02e-3 9.12e-8
B2B [25] 94m2s 499,758 11.74e-3 27.31e-8
DeepONet [25] 4m10s 504,926 115.14e-3 2588.68e-8
FNO2D [34] 31m22s 1,188,353 5.43e-3 5.19e-8
RFM-C2C 5s+3m17s 492,864 6.47e-3 8.68e-8
FEM-C2C 7m27s+4m51s 504,979 6.92e-3 9.03e-8
4.3 Poisson equation in the complex domain
In this example, we consider a two-dimensional Poisson equation on the complex domain $\Omega$ defined as
$\Delta u(x)$ $=f(x),\quad x\in\Omega,$ (29)
$u(x)$ $=0,\quad x\in\partial\Omega,$
where the domain $\Omega$ is a rectangle with three interior holes of different shapes. The objective is to learn the solution operator
$G:f(x)\mapsto u(x).$
The forcing term $f(x)$ is sampled from the function space
$f(x)=\sum_{i=0}^{4}\sum_{j=0}^{4}\alpha_{ij}\sin\!\left([i\pi,j\pi]\cdot x^{\top}\right),$
where the coefficients $\alpha_{ij}$ are independently drawn from the uniform distribution on [1,1][-1,1]. The reference solutions are generated using the finite element method (FEM) on a mesh with 1780 nodes as shown in Figure 9 (a).
In this experiment, we employ both FEM-C2C and RFM-C2C methods. The training set consists of 1600 samples and the testing set consists of 400 samples. The basis numbers for encoder and decoder are set to $(m_{1},m_{2})=(533,1586)$ and $(m_{1},m_{2})=(512,1600)$, respectively. We adopt a piecewise exponential learning rate decay: The learning rate for the FEM-C2C method starts from 1e51e-5 and is multiplied by 0.9 every 1000 steps for a total of 30000 training steps. The initial rates are 1e31e-3, 1e41e-4, and 1e51e-5 for three consecutive 10000-step segments, with a decay factor of 0.9 applied every 1000 steps within each segment for RFM-C2C. The FEM basis meshes for encoder and decoder are shown in Figure 9 (b) and (c). The mean relative $L^{2}$ errors on the test set is 4.13e34.13e-3 for FEM-C2C and 7.69e37.69e-3 for RFM-C2C method. The worst case is shown in Figure 10 and Figure 11, respectively.
Relation between Basis Approximation and FB-C2C Error. In this experiment, we aim to verify that the accuracy of the proposed FB-C2C method is influenced by the approximation capability of the output basis system. As illustrated in Figure 12 (a) and (b), the red dashed line represents the approximation error of the output basis system to the output function space, while the blue solid line represents the accuracy of the FB-C2C method in solving the problem. It can be observed that the blue curve always lies above the red dashed curve, indicating that the accuracy of FB-C2C is bounded by the approximation accuracy of the output basis system (i.e., the error cannot be smaller than the approximation error). In this experiment, increasing the number of output basis functions $m_{2}$ generally improves the approximation capability of the output basis system. This process is mathematically formulated in the previous inequality (26). Consequently, as $m_{2}$ increases, both the approximation error and the FB-C2C error decrease. The fact that the two curves nearly coincide further demonstrates that FB-C2C is sufficiently trained in this setting.
(a) Sample grid points for the functions.
(b) FEM mesh for encoder.
(c) FEM mesh for decoder.
Figure 9: 2D Poisson equation: Data representation and FEM mesh discretization. (a) Sample grid points for the functions. (b) FEM mesh for the encoder. (c) FEM mesh for the decoder.
Figure 10: 2D Poisson equation: Worst-case sample from the complex domain Poisson dataset solved by FEM-C2C. (a) Pointwise comparison between the prediction and the ground truth. (b) Exact solution. (c) Predicted solution.
Figure 11: 2D Poisson equation: Worst-case sample from the complex domain Poisson dataset solved by RFM-C2C. (a) Pointwise comparison between the prediction and the ground truth. (b) Exact solution. (c) Predicted solution.
(a)FEM basis.
(b) RFM basis.
Figure 12: 2D Poisson equation: Prediction error of the learned operator versus the best possible projection error in the fixed basis space. The C2C error (blue) is bounded below by the approximation error of the output basis system (red), and both decrease as the number of output bases $m_{2}$ increases. (a) FEM basis. (b) RFM basis.
4.4 Elastic plate
In this example, we demonstrate the capability of the proposed method to approximate solution operators for problems with coupled multi-field outputs. This investigation concerns a thin rectangular plate subjected to in-plane loading, described by the two-dimensional plane stress elasticity equations under the assumption of negligible body forces. The computational domain is defined as $\Omega=(0,1)\times(0,1)$ with a thin rectangular region removed from the interior, representing a cut-out in the plate. The equilibrium equations are
$\begin{array}[]{cc}\nabla\cdot\mathbf{\sigma}+\mathbf{b}(\mathbf{x})=\mathbf{0},\quad\mathbf{x}=(x,y)\in\Omega,\end{array}$ (30)
where $\mathbf{\sigma}=\left[\begin{matrix}\sigma_{xx}&\tau_{xy}\\ \tau_{xy}&\sigma_{yy}\end{matrix}\right]$ denotes the Cauchy stress symmetric tensor ($\tau_{xy}=\tau_{yx}$), $\mathbf{b}(\mathbf{x})=\mathbf{0}$ is the body force. The material parameters are the Young’s modulus $E=300\times 10^{5}$ and Poisson’s ratio $\nu=0.3$. Under plane stress conditions, the constitutive relation between stress and displacement gradients is given by
$\begin{bmatrix}\sigma_{xx}\\ \sigma_{yy}\\ \tau_{xy}\end{bmatrix}=\frac{E}{1-\nu^{2}}\begin{bmatrix}1&\nu&0\\ \nu&1&0\\ 0&0&\frac{1-\nu}{2}\end{bmatrix}\begin{bmatrix}\epsilon_{xx}\\ \epsilon_{yy}\\ \gamma_{xy}\end{bmatrix}$. (31)
The kinematic equations relate the displacements to the strains as
$\epsilon_{xx}=\frac{\partial u}{\partial x},\quad\epsilon_{yy}=\frac{\partial v}{\partial y},\quad\gamma_{xy}=\frac{\partial u}{\partial y}+\frac{\partial v}{\partial x}.$ (32)
The boundary conditions consist of fixed displacements on the left edge of the plate and applied traction on the right edge, expressed as
[…]
where $\mathbf{n}$ is the unit outward normal vector on the boundary, $\mathbf{f}(\mathbf{x})=[0,f(1,y)]^{T}$ is the boundary load applied at the edge $x=1$, and $u(\mathbf{x})$ and $v(\mathbf{x})$ represent the displacements in the $x$- and $y$-directions, respectively. The boundary loading $f(1,y)$ applied to the right edge of the plate is modeled as a Gaussian random field. Our objective is to learn the map:
$G:f(1,y)\mapsto[u(\mathbf{x}),v(\mathbf{x})].$
The dataset for this problem is adopted from [25]. The input grid points are uniformly sampled over the domain with $101$ points, while the output grid contains $1048$ points, which is shown in Figure 13 (a). In total, $1850$ samples are used for training and $100$ samples are reserved for testing.
We now examine an operator with a multi-function output (specifically, two functions), whereas previous cases involved single-output operators. The following section details the two approaches pursued for this scenario: the scalar-valued and vector-valued basis methods.
Scalar-Valued basis functions for Multi-Output Function Approximation. We use scalar-valued basis functions $\{\psi_{i}(\mathbf{x})\}_{i=1}^{m_{2}}$ to approximate $u(\mathbf{x})$ and $v(\mathbf{x})$, i.e.,
$u(\mathbf{x})\approx\sum_{j=1}^{m_{2}}b_{j}^{u}\,{\psi}_{j}(\mathbf{x}),\quad v(\mathbf{x})\approx\sum_{j=1}^{m_{2}}b_{j}^{v}\psi_{j}(\mathbf{x}),$ (34)
where the corresponding output coefficients $\mathbf{b}=[b_{1}^{u},b_{2}^{u},..,b_{m_{2}}^{u},b_{1}^{v},b_{2}^{v},...,b_{m_{2}}^{v}]\in\mathbb{R}^{m_{2}\times 2}$ associated with the scalar-valued basis space $V_{\text{out}}=\text{span}\{\psi_{i}(\mathbf{x})\}_{i=1}^{m_{2}}$.
Vector-Valued basis functions for Multi-Output Function Approximation. In this example, we denote vector-valued basis space
$V_{\text{out}}=\text{span}\{\bm{\psi}_{i}(\mathbf{x})\}_{i=1}^{m_{2}}=\text{span}\{[\psi_{i}^{u}(\mathbf{x}),\psi_{i}^{v}(\mathbf{x})]\}_{i=1}^{m_{2}},$
which are employed to approximate the multi-output function space $\mathcal{Y}$. A given output multiple functions $\bm{u}(x)=[u(x),v(x)]\in\mathcal{Y}$ is represented in terms of the vector-valued basis functions, i.e.,
$\bm{u}(\mathbf{x})=[u(\mathbf{x}),v(\mathbf{x})]\approx\sum_{j=1}^{m_{2}}b_{j}\,\bm{\psi}_{j}(\mathbf{x})=\sum_{j=1}^{m_{2}}b_{j}\,[\psi_{j}^{u}(\mathbf{x}),\psi_{j}^{v}(\mathbf{x})],$ (35)
where $\mathbf{b}=[b_{1},b_{2},...,b_{m_{2}}]\in\mathbb{R}^{m_{2}}$ are the expansion coefficients associated with multiple output functions $\bm{u}=[u,v]$.
Scalar-Valued vs. Vector-Valued Basis Functions: Efficiency and Accuracy. Both methods use $m_{1}=64$ input basis functions, a count precisely set in one dimension. Their output dimensions differ: in the 2D scalar-valued case, RFM attains $m_{2}=512$ while FEM yields $m_{2}=540$, as the latter cannot precisely mesh 512 triangles; in the vector-valued case, randomness generates a different basis, but these $m_{2}$ values are preserved. When we generate scalar valued basis functions, we just set $\psi_{i}=\psi_{i}^{u}$. Two different meshes for the FEM basis generation are shown in Figure 13 (b) and Figure 13 (c).
Some experimental results are shown in Table 5. As the vector-valued basis requires only half the output coefficients in this example, it contains significantly fewer trainable parameters, resulting in faster convergence and enhanced training efficiency. The relative error curves in Figure 14 (RFM) show similar accuracy for scalar-valued and vector-valued bases. Conversely, Figure 15 (FEM) reveals a significant accuracy gap, with the vector-valued basis performing much worse than the scalar-valued one. Therefore, the vector-valued basis provides superior training efficiency over the scalar-valued basis and, as realized in the RFM-C2C method, markedly better accuracy than the vector-valued FEM-C2C method.
Table 5: Elastic plate: The comparison of C2C framework and P2C framework in scalar-valued and vector-valued RFM and FEM basis spaces. The cost time for encoder means the time required to generate the basis functions and project input functions to the corresponding basis space. The training time refers to the time required to train the neural operator.
Cost time Training time Test relative $L^{2}$ error Neural network structure
Basis Encoder C2C P2C C2C P2C C2C P2C
Scalar RFM 0.82s 43.40s 48.11s 8.47e-3 141.56e-3 [64,512,1024] [101,512,1024]
Scalar FEM 11.75s 49.01s 52.80s 1.68e-3 1.93e-3 [64,512,1080] [101,512,1080]
Vector RFM 0.82s 42.11s 64.86s 7.81e-3 140.81e-3 [64,512,512] [101,512,512]
Vector FEM 11.75s 40.22s 42.84s 626.44e-3 626.43e-3 [64,512,540] [101,512,540]
(a) Sample grid points for output functions.
(b) $\psi^{v}$ FEM basis mesh.
(c) $\psi^{u}$ FEM basis mesh.
Figure 13: Elastic plate: Grid points and two FEM basis meshes for the output functions.
(a) Scalar-Valued RFM basis.
(b) Vector-Valued RFM basis.
Figure 14: Elastic plate: Training curve comparison between the scalar-valued and vector-valued RFM basis. The number of grid points is set to $n_{1}=101$ and $n_{2}=1048$, and the number of basis functions is $m_{1}=64$ and $m_{2}=512$.
(a) Scalar-Valued FEM basis.
(b) Vector-Valued FEM basis.
Figure 15: Elastic plate: Training curve comparison between the scalar-valued and vector-valued FEM basis. The number of grid points is set to $n_{1}=101$ and $n_{2}=1048$, and the number of basis functions is $m_{1}=64$ and $m_{2}=540$.
4.5 Darcy Flow in an L-shaped domain
We investigate Darcy flow in a two-dimensional L-shaped region to validate our approach in handling multi-input scenarios. A distinctive feature of this example is that it involves a mapping from two functions to a single function, thereby highlighting the broad applicability of our method. The problem is formulated as
[…]
$u(\mathbf{x})$ $=g(\mathbf{x}),\quad\mathbf{x}\in\partial\Omega,$
where $k(\mathbf{x})$, $u(\mathbf{x})$, and $f(\mathbf{x})$ denote the permeability field, hydraulic head, and source term, respectively, all exhibiting spatial variability. The dataset is from [25]. Its training set consists of 51000 samples and the testing set consists of 7000 samples. Figure 16 shows the discretization of input and output functions over sampling grids of 736 and 450 points, respectively. In this case we aim to approximate the map:
$G:[k(\mathbf{x}),f(\mathbf{x})]\mapsto u(\mathbf{x}).$
This example departs from the preceding cases by considering an operator that maps two input functions to a single output. To enhance both training efficiency and accuracy, we leverage vector-valued basis functions, specifically employing the RFM type motivated by our earlier findings.
Vector-Valued Basis Functions for Multi-Input Function Approximation. For the approximation of a single function, as considered in previous numerical experiments, we may use a linear space spanned by a set of scalar-valued basis functions. However, for the joint approximation of a family of functions consisting of multiple components, it is therefore better to work within a vector space spanned by vector-valued basis functions, as it leads to superior training efficiency.
In this example, we define the input basis system as follows.
$V_{\text{in}}=\text{span}\{\bm{\phi}_{i}(\mathbf{x})\}_{i=1}^{m_{1}}=\text{span}\{[\phi_{i}^{f}(\mathbf{x}),\phi_{i}^{k}(\mathbf{x})]\}_{i=1}^{m_{1}},$
which is used to approximate the input function space $\mathcal{X}$. A given input pair of functions $\mathbf{f}(\mathbf{x})=[f(\mathbf{x}),k(\mathbf{x})]\in\mathcal{X}$ can then be represented in terms of these vector-valued basis functions:
$\mathbf{f}(\mathbf{x})=[f(\mathbf{x}),k(\mathbf{x})]\approx\sum_{j=1}^{m_{1}}a_{j}\,\bm{\phi}_{j}(\mathbf{x})=\sum_{j=1}^{m_{1}}a_{j}\,[\phi_{j}^{f}(\mathbf{x}),\phi_{j}^{k}(\mathbf{x})],$ (37)
where $\{a_{j}\}_{j=1}^{m_{1}}$ are the corresponding expansion coefficients.
The coefficients $\{a_{j}\}$ are obtained by solving the following vector-valued regularized least-squares problem:
$\min_{\mathbf{a}\in\mathbb{R}^{m_{1}}}\;\frac{1}{n}\sum_{i=1}^{n}\left\|\mathbf{f}(\mathbf{x}_{i})-\sum_{j=1}^{m_{1}}a_{j}\,\bm{\phi}_{j}(\mathbf{x}_{i})\right\|_{2}^{2}+\lambda\|\mathbf{a}\|_{2}^{2},$ (38)
where $\{\mathbf{x}_{i}\}_{i=1}^{n}$ are the sampling points in the domain, and the second term provides $L^{2}$-regularization to stabilize the coefficient estimation.
Generalization Performance of RFM-C2C vs. RFM-P2C. In this example, we compare the generalization error of RFM-C2C and RFM-P2C methods by examining the effect of varying the number of training samples on the test results. We adopt a piecewise exponential learning rate decay: The learning rate for both methods starts from 1e31e-3 and is multiplied by 0.9 every 1000 steps for a total of 10000 training steps. The generalization capabilities of both methods under fixed RFM basis spaces are evaluated using training sets of 1000 samples, 5100 samples, and the complete dataset. As shown in Figure 17, the C2C method consistently achieves superior generalization performance and higher accuracy compared to the P2C method across all three training sets. This advantage becomes increasingly evident as the amount of training data decreases, demonstrating the robustness of the C2C method in data-scarce scenarios. Furthermore, the worst-case relative error from the L-shaped Darcy flow test set, obtained with full training data using the C2C method, is illustrated in Figure 17.
Table 6: Darcy flow in an L-shaped domain: The comparison of vector-valued C2C framework and P2C framework training on different train sets. The training time refers to the time required to train the neural operator.
Training time Test relative $L^{2}$ error Neural network structure
Data number C2C P2C C2C P2C C2C P2C
1000 19.11s 20.59s 253.92e-3 261.67e-3 [512,256,256,1024] [1472,256,256,1024]
5100 20.96s 24.01s 67.04e-3 96.21e-3 [512,256,256,1024] [1472,256,256,1024]
51000 156.48s 186.67s 38.55e-3 42.09e-3 [512,256,256,1024] [1472,256,256,1024]
(a) Distribution of input function sample points.
(b) Distribution of output function sample points.
Figure 16: Darcy flow in an L-shaped domain: Sample grid points for functions.
(a) 1000 training data
(b) 5100 training data
(c) 51000 training data
Figure 17: Darcy flow in an L-shaped domain: Training curve comparison between the RFM-C2C and RFM-P2C methods in three training datasets of varying scales. All three test datasets are identically sized, each comprising 7000 instances. The number of grid points is set to $n_{1}=736\times 2$ and $n_{2}=450$, and the number of RFM basis functions is $m_{1}=512$ and $m_{2}=1024$.
Figure 18: Darcy flow in an L-shaped domain: Worst-case sample from the L-shaped Darcy flow dataset with 51000 training data. (a) Pointwise comparison between the prediction and the ground truth. (b) Exact solution. (c) Predicted solution.
4.6 KdV-Burgers Equation
To illustrate the capacity of the proposed framework for solving nonlinear time-dependent partial differential equations, we investigate the KdV-Burgers equation with periodic boundary conditions:
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}$ $=\nu\frac{\partial^{2}u}{\partial x^{2}}-\mathcal{D}\frac{\partial^{3}u}{\partial x^{3}},$ $x\in(0,1),\ t\in(0,1),$ (39)
$u(0,t)$ $=u(1,t),$ $t\in[0,1]$
$\frac{\partial u}{\partial x}(0,t)$ $=\frac{\partial u}{\partial x}(1,t),$ $t\in[0,1]$
$u(x,0)$ $=f(x),$ $x\in[0,1]$
where the viscosity coefficient $\nu=0.01$ and the dispersion coefficient $\mathcal{D}=0.001$. The initial condition $f(x)$ is drawn from a Gaussian random field $\mathcal{N}\left(0,25^{2}(-\Delta+25I)^{-4}\right)$, consistent with the boundary constraints. The referenced solutions are generated using a pseudo-spectral approach discretized over a uniform spatial grid comprising 8192 nodes, integrated in time via a forward Euler scheme with a fine temporal resolution, as outlined in [24]. The resolutions of the input function and output function are 101 and $101\times 101$ respectively. The operator maps the initial condition $f(x)$ to the spatiotemporal solution $u(x,t)$:
$G:f(x)\mapsto u(x,t).$
Fully Coupled Spatiotemporal Basis Functions for Time-Dependent Solution Approximation. To effectively capture the spatiotemporal dynamics of the KdV-Burgers equation, we employ non-separable spatiotemporal basis functions according to RFM basis system (9). Specifically, we define the output basis function as
$\phi(t,x)=\sum_{j=1}^{M}a_{j}\sigma(w_{j}t+k_{j}x+b_{j}),$ (40)
where the inner parameters $\{w_{j},k_{j},b_{j}\}_{j=1}^{M}$ are randomly initialized. The initial learning rates are 1e31e-3, 1e41e-4, and 1e51e-5 for three consecutive 10000-step segments, with a decay factor of 0.9 applied every 600 steps within each segment. The mean relative $L^{2}$ error is 3.29e-3 on the test dataset by RFM-C2C, which verifies the effectiveness of the proposed method for time-dependent PDEs. The worst-case scenario for the KdV-Burgers equation is illustrated in Figure 19.
Figure 19: KdV-Burgers equation: Worst-case sample from the KdV-Burgres equation dataset solved by RFM-C2C. (a) Pointwise comparison between the prediction and the ground truth. (b) Exact solution. (c) Predicted solution.
4.7 2D Helmholtz Equation
In contrast to the previous example with strong solutions, the proposed method can learn mappings to weak solution spaces. We consider the Helmholtz equation defined on a 2 dimensional rectangular domain $\Omega=\{(x,y)\in\mathbb{R}^{2}\mid x\in(-1,1),y\in(0,1)\}$ with parameter $\lambda=1$:
$\begin{cases}\Delta u+\lambda u=f(x,y),\quad(x,y)\in\Omega\\ u|_{\partial\Omega}=\alpha\sinh(|x-\beta|)\cos(y^{2}),\end{cases}$ (41)
where $f(x,y)=\alpha\sinh(|x-\beta|)\left((1+\lambda-4y^{2})\cos(y^{2})-2\sin(y^{2})\right)$, $\alpha$ and $\beta$ are two constants. This equation does not admit a strong solution due to the lack of differentiability at $x=0$, though a weak solution exists. The weak solution is given by $u(x,y)=\alpha\sinh(|x-\beta|)\cos(y^{2})$. To construct the dataset, we choose $50$ random values of $\alpha\in U([1,2])$ and $50$ random values of $\beta\in U([-0.5,0.5])$. Thus, the total data number is 2500. We use 2000 samples for training and 500 samples for testing. The resolutions of the input function and output function are both uniformly discretized as $101\times 101$. We learn the operator that maps forcing term $f(x,y)$ to the weak solution $u(x,y)$:
$G:f(x,y)\mapsto u(x,y).$
Weak Solution Operator Learning. Even though solutions may not be smooth, the proposed method can still learn the weak solution operator $G$ by RFM-C2C (FEM-C2C). The output basis functions are constructed as in (9). The number of basis functions is $m_{1}=256$ $(m_{1}=273)$ and $m_{2}=512$ $(m_{2}=558)$. The RFM basis functions are smooth but can approximate non-smooth solutions in the weak sense. The learning rate for both methods starts from 1e31e-3 and is multiplied by 0.9 every 200 steps for a total of 10000 training steps. The main error occurs near the non-differentiable line. To show clearly, we plot the predicted solution and boundary condition comparison on the boundary $y=0$ and $y=1$. The worst-case scenario for the 2D Helmholtz equation is illustrated in Figure 20 (Figure 21). The mean relative $L^{2}$ error is 6.51e-3 (2.13e-3) in the test dataset by RFM-C2C (FEM-C2C), which verifies the effectiveness of the proposed method for weak solution learning.
Figure 20: 2D Helmholtz equation: The worst-case predicted solution obtained by RFM-C2C method. (a) Pointwise comparison between the prediction and the ground truth. (b) Exact solution. (c) Predicted solution. (d) Predicted solution and boundary condition on $y=0$. (e) Predicted solution and boundary condition on $y=1$.
Figure 21: 2D Helmholtz equation: The worst-case predicted solution obtained by FEM-C2C method. (a) Pointwise comparison between the prediction and the ground truth. (b) Exact solution. (c) Predicted solution. (d) Predicted solution and boundary condition on $y=0$. (e) Predicted solution and boundary condition on $y=1$.
4.8 High Dimension Poisson Equation
In this example, we consider operator learning for high-dimensional functions. In this setting, if the finite element method is employed to construct the basis functions, one inevitably encounters the curse of dimensionality, as the number of basis functions required for an accurate approximation grows exponentially with the spatial dimension. Therefore, in such cases, adopting the RFM-C2C framework provides a more natural and efficient alternative. Also, in such cases, the number of grid points typically grows exponentially with the dimension, making conventional point-to-coefficient (P2C) frameworks impractical. By employing random feature models (RFM) as the basis functions, our approach effectively alleviates this issue and reduces the complexity of the neural operator network. To illustrate the capability of our method in addressing high-dimensional problems, we present the following example on the 10 dimensional ($d=10$) Poisson equation:
$-\Delta u=$ $f\quad\text{in }\Omega,$ (42)
$u=$ $\alpha\sum_{i=1}^{d}\sin\!\left(\tfrac{\pi}{2}x_{i}\right)$ $\quad\text{on }\partial\Omega,$
where $\Omega=(-1,1)^{d}$ and $f(\mathbf{x})=\alpha\frac{\pi^{2}}{4}\sum_{i=1}^{d}\sin\left(\frac{\pi}{2}x_{i}\right)$, which admits the exact solution
$u(\mathbf{x})=\alpha\sum_{i=1}^{d}\sin\left(\frac{\pi}{2}x_{i}\right).$
Here the objective is to learn the map
$G:f(\bm{x})\mapsto u(\bm{x}).$
To construct the training dataset, we uniformly sample $31^{10}$ points from the domain $1,1]^{10}$ and choose $8$ random values of $a\in[-1,1]$.
The solutions are then evaluated for different values of $a$. In particular, we designate as Generation 1 (Gen1) the case of $2$ test values with $a\in[-1,1]$ that are not included in the training set. Furthermore, we perform extrapolation tests on $10$ test values of $a\in[1,2]$, which lie outside the training range and are designated as Generation 2 (Gen2). When training the RFM to obtain the coefficients, we employ the stochastic gradient descent (SGD) method. At each training step, we use a batch size of $1e5$, i.e., we randomly select $10^{5}$ points from the total $31^{10}$ grid points to compute the RFM coefficients. While the P2C approach struggles with the immense scale of $31^{10}$, the C2C method effectively addresses this by reducing the input dimension from $31^{10}$ to a manageable size $m_{1}$.
In this example, we adopt the random feature model as the basis functions and use the same basis for both the input and output spaces, i.e., $m_{1}=m_{2}$. We compare two settings with $m_{1}=m_{2}=128$ and $m_{1}=m_{2}=8192$, respectively. The learning rate for both methods starts from 1e31e-3 and is multiplied by 0.9 every 200 steps for a total of 10000 training steps. From the results in Table 7, we observe that our RFM-based C2C method is able to effectively handle high-dimensional problems while achieving strong generalization. Using a smaller number of basis functions can accelerate training, though at the expense of a slight loss in accuracy. The training and testing curves are presented in Figure 22, while the worst-case sample for the Gen1 when $m_{1}=m_{2}=8192$ is illustrated in Figure 23.
Table 7: High-dimensional Poisson equation: Mean relative $L^{2}$ errors between the predicted and exact solutions $u_{i}(\mathbf{x})$ for different numbers of basis functions. The training time refers to the time required to train the neural operator.
$m_{1}=m_{2}=128$ $m_{1}=m_{2}=8192$
Dimension $d$ Gen1 Gen2 Training Time Gen1 Gen2 Trianing Time
10 $1.69\mathrm{e\text{-}02}$ $3.00\mathrm{e\text{-}02}$ 1208.5s $6.47\mathrm{e\text{-}03}$ $1.27\mathrm{e\text{-}02}$ 76663.5s
Figure 22: High-Dimensional Poisson equation: Training curves comparison for different numbers of RFM basis functions.
Figure 23: High-Dimensional Poisson equation: Worst-case sample from the high-dimensional Poisson Gen1 dataset when $m_{1}=m_{2}=8192$. Visualization of $u(\mathbf{y})$ along the cross-section $(y_{3},y_{4})$ with all other coordinates fixed as $y_{i}=0$ for $i\neq 3,4$. (a) Pointwise comparison between the prediction and the ground truth. (b) Exact solution. (c) Predicted solution.
5 Inverse Boundary Problem for the Stokes System
Forward Problem
Let $\Omega_{0}\subset\mathbb{R}^{2}$ be a fixed computational domain and $\mathcal{O}\subset\Omega_{0}$ denote an obstacle, possibly consisting of multiple connected components. The fluid domain is defined by $\Omega=\Omega_{0}\setminus\mathcal{O},\Gamma=\partial\mathcal{O}.$ We consider the steady incompressible Stokes equations in $\Omega$,
$\begin{cases}-\mu\Delta\mathbf{u}+\nabla p=\mathbf{0},&\mathbf{x}\in\Omega,\\ \nabla\cdot\mathbf{u}=0,&\mathbf{x}\in\Omega,\\ \mathbf{u}=\mathbf{0},&\mathbf{x}\in\Gamma\quad\text{(no-slip)},\\ \mathbf{u}=\mathbf{U}_{\infty},&\mathbf{x}\in\partial\Omega_{0}.\end{cases}$ (43)
Here $\mathbf{u}=(u_{x},u_{y})^{\top}$ denotes the velocity field, $p$ is the pressure, $\mu>0$ is the dynamic viscosity, and $\mathbf{U}_{\infty}$ is a prescribed far-field velocity. The obstacle geometry is represented by a signed distance function (SDF) $\phi:\Omega_{0}\to\mathbb{R}$, defined by
$\phi(\mathbf{x})=\begin{cases}-\mathrm{dist}(\mathbf{x},\Gamma),&\mathbf{x}\in\mathcal{O},\\ \phantom{-}\mathrm{dist}(\mathbf{x},\Gamma),&\mathbf{x}\in\Omega,\end{cases}\qquad\Gamma=\{\mathbf{x}\in\Omega_{0}:\phi(\mathbf{x})=0\}.$ (44)
For a given $\phi$, the forward problem (43) admits a unique velocity solution $\mathbf{u}(\mathbf{x};\phi)$. The solution can be obtained by reformulating the problem as a boundary integral equation. Here, we consider a inverse problem setting.
Inverse problem: sparse velocities-to-geometry
In the inverse setting, the full velocity field $\mathbf{u}(\mathbf{x})$ is not available. Instead, only sparse velocity measurements at a limited number of sensor locations $\{\mathbf{x}_{i}\}_{i=1}^{N_{s}}\subset\Omega_{0}$ are observed:
$\mathcal{Y}=\left\{\bigl(\mathbf{x}_{i},\mathbf{u}(\mathbf{x}_{i})\bigr)\right\}_{i=1}^{N_{s}},\qquad\mathbf{u}(\mathbf{x}_{i})=(u_{x}(\mathbf{x}_{i}),u_{y}(\mathbf{x}_{i})).$ (45)
These sensor locations may be confined to a few spatial windows downstream of the obstacle. In the numerical experiments, we consider a two-dimensional computational domain $\Omega_{0}=[-3,3]\times[-3,3].$ The obstacle $\mathcal{O}$ is assumed to be fully contained within the subdomain $1.5,1.5]\times[-1.5,1.5]\subset\Omega_{0}.$ Velocity measurements are available only in several prescribed observation windows located downstream of the obstacle. Specifically, the sensor locations are restricted to the union of the following regions: $1.5,3.0]\times[-3.0,3.0],[0.5,3.0]\times[-3.0,-2.0],[0.5,3.0]\times[2.0,3.0].$ The dataset comprises 2,500 training samples and 500 testing samples, featuring diverse obstacle geometries such as circular, elliptical, and star-shaped shapes, with one to three obstacles of varying sizes and locations.
The inverse problem consists of recovering the obstacle geometry (equivalently, the signed distance function $\phi$) from the sparse velocity observations:
$\text{Given }\mathcal{Y},\quad\text{find }\phi\text{ such that }\mathbf{u}(\mathbf{x}_{i};\phi)\approx\mathbf{u}(\mathbf{x}_{i}),\quad i=1,\dots,N_{s}.$ (46)
The velocity field is encoded using a random feature model (RFM), whereas the obstacle geometry is parameterized by a fixed random radial basis function (RBF) expansion,
$\phi(\mathbf{x})=\sum_{m=1}^{M}c_{m}\,\psi_{m}(\mathbf{x}),\qquad\psi_{m}(\mathbf{x})=\exp\!\left(-\frac{\|\mathbf{x}-\bm{\mu}_{m}\|^{2}}{2\sigma_{m}^{2}}\right),$ (47)
where the centers $\{\bm{\mu}_{m}\}$ and widths $\{\sigma_{m}\}$ are randomly sampled and fixed. The results are shown below:
Figure 24: Inverse boundary problem. Reconstruction results for different obstacle geometries. The red dashed curves denote the ground-truth boundaries, while the black solid curves represent the recovered obstacles. Black dots indicate the sensor locations, with a total of $N_{s}=100$ velocity measurement points. Overall, the relative $L_{2}$ error is 5.562e25.562e-2.
6 Conclusion
We propose a coefficient-to-coefficient (C2C) operator learning framework built upon fixed input and output basis spaces, termed the Fixed-Basis Coefficient-to-Coefficient Operator Network (FB-C2CNet). In contrast to conventional neural operator frameworks that rely on pointwise evaluations on spatial grids, our framework operates directly on the coefficients associated with a set of pre-selected basis functions. This coefficient-based representation significantly reduces the input dimensionality and the number of trainable parameters, thereby mitigating training difficulty and improving computational efficiency.
Within this framework, we consider several representative choices of basis functions, including Random Feature Model (RFM) bases, Finite Element Method (FEM) bases and radial basis function (RBF). We show that the effective rank of the encoded coefficient matrix plays a crucial role in determining the generalization error of the learned operator. In particular, a larger effective rank of the coefficient representation consistently leads to improved generalization performance. In addition, we provide a theoretical analysis demonstrating that the accuracy of FB-C2CNet is limited by the projection error associated with the prescribed output basis space $V_{\mathrm{out}}$.
We conduct extensive numerical experiments to evaluate the proposed FB-C2CNet across a wide spectrum of operator learning tasks, including linear and nonlinear operators, regular and complex geometries, one-, two-, and high-dimensional problems, as well as inverse settings. The results demonstrate that FB-C2CNet achieves Competitive predictive fidelity with remarkable computational efficiency. Notably, compared to Point-to-Coefficient (P2C) methods, the proposed framework exhibits superior generalization capabilities, particularly in challenging scenarios.
It is noteworthy that by prescribing the basis functions a priori, the proposed framework substantially reduces the number of trainable parameters while maintaining high-fidelity predictive performance. Consequently, FB-C2CNet demonstrates significantly reduced training overhead compared to conventional neural operator models. Furthermore, this fixed-basis formulation establishes a natural interface with classical numerical methods, enabling the framework to leverage established discretization techniques and facilitating the development of hybrid algorithms that combine data-driven efficiency with numerical reliability.
Acknowledgments
The work of Y.X and W.H.Z was supported in part by the National Key Research and Development Program of China (No. 2025YFA1016800) and the Project of Hetao Shenzhen-HKUST Innovation Cooperation Zone HZQB-KCZYB-2020083.
Appendix
Experiments Details
Unless otherwise stated, the learning rate follows the scheduling scheme illustrated in Figure 25.
Figure 25: Learning Rate Schedule. The initial rates are 1e21e-2, 1e31e-3, and 1e41e-4 for three consecutive 10000-step segments, with a decay factor of 0.9 applied every 200 steps within each segment
In our method, the quantities shown in Table 8 are primarily considered. Some parameters for each dataset and the RFM basis systems are shown in Table 9.
Table 8: Key quantities and notation.
Quantity Notation
Number of training data $N_{\text{train}}$
Number of testing data $N_{\text{test}}$
Datasets Number of input grid points $n_{1}$
Number of output grid points $n_{2}$
Number of input basis functions $m_{1}$
Basis systems Number of output basis functions $m_{2}$
Initial Variance $R_{m}$
Random Feature Model Number of PoU $M_{p}$
Cut-off cutcut
Table 9: The parameters for each dataset and input (output) RFM basis systems.
$N_{\text{train}}$ $N_{\text{test}}$ $n_{1}$ $n_{2}$ $m_{1}$ $m_{2}$ $R_{m}$ $M_{p}$ cutcut
1D Darcy Flow 1000 800 2000 2000 128 128 3 (3) [16] ([16]) 1e-1
1D Darcy Flow 1000 800 100 100 128 128 3 (3) [8] ([4]) 1e-2
2D Darcy Flow (regular) 2400 600 19881 19881 512 2048 0.3 (0.3) [8,8] ([8,8]) 1e-1
2D Poisson Equation 1600 400 1780 1780 512 1600 3 (3) [2,2] ([4,4]) 1e-2
Elastic Plate Equation 1850 100 101 $1048\times 2$ 64 512 3 (3) [4] ([2,2]) 1e-3
2D Darcy Flow (Lshaped) $51000$ 7000 $736\times 2$ 450 512 1024 1 (3) [8,8] ([8,8]) 1e-2
2D Darcy Flow (Lshaped) $5100$ 7000 $736\times 2$ 450 512 1024 1 (3) [8,8] ([8,8]) 1e-2
2D Darcy Flow (Lshaped) $1000$ 7000 $736\times 2$ 450 512 1024 1 (3) [8,8] ([8,8]) 1e-2
KdV-Burgers Equation 1600 400 101 19881 32 1024 3 (1) 4 ([8,8]) 1e-2
2D Helmholtz Equation 2000 500 19881 19881 256 512 3 (1) [4,4] ([4,4]) 1e-8
10D Poisson Equation 8 2 $31^{10}$ $31^{10}$ 128 128 0.3(0.3) [1,1,1,1,1,1,1,1,1,1] 1e-1
10D Poisson Equation 8 2 $31^{10}$ $31^{10}$ 8192 8192 0.3 (0.3) [1,2,1,1,2,1,2,2,1,2] 1e-1
Inverse Stokes system 2500 500 $100$ $128^{2}$ $64$ $1024$ 1 ([0.03,2.4]) [2,2] (-) 1e-2
Experiments Results
The performance of our method is listed in Table 10 and Table 11. The dash (-) in Table 10 is used to represent cases where no experiment was carried out. The multiplication sign ($\times$) indicates layer repetition: 512×5 denotes five consecutive layers, each having 512 neurons.
Table 10: Mean relative $L^{2}$ error of the testing data.
Scalar FEM-C2C Scalar RFM-C2C Vector FEM-C2C Vector RFM-C2C
1D Darcy Flow ($n_{1}=n_{2}=2000$) 13.67e-3 3.87e-3 - -
2D Darcy Flow (Regular) 6.92e-3 6.47e-3 - -
2D Poisson Equation 4.13e-3 7.69e-3 - -
Elastic Plate Equation 1.68e-3 8.47e-3 624.43e-3 7.81e-3
2D Darcy Flow (Lshaped) - - - 38.55e-3
KdV-Burgers Equation - 3.29e-3 - -
2D Helmholtz Equation 2.13e-3 6.51e-3 - -
10D Poisson Equation - 6.48e-3 - -
Inverse Stokes system - 5.56e-2 - -
Table 11: Performance comparison for different benchmarks. Performance are measured with relative $L^{2}$ errors ($RL^{2}E$) and mean squared errors (MSEMSE). The structure of the neural networks used in each experiment is detailed in the “Net” column of the table. Net for BasisONet and B2B includes input basis generator, transformation network and output basis generator. Net for DeepONet means the branch network++ trunk network ++ bias term for the dot product result.
2D Darcy Flow Time Params $RL^{2}E$ MSEMSE Net
[…]
FNO2D 31m22s 1,188,353 5.43e-3 5.19e-8 modes = 12, width = 32, epochs=500
[…]
2D Poisson time params $RL^{2}E$ MSEMSE Net
[…]
Elastic Plate time params $RL^{2}E$ MSEMSE Net
[…]
LShaped Darcy time params $RL^{2}E$ MSEMSE Net
[…]
KdV-Burgers time params $RL^{2}E$ MSEMSE Net
[…]
2D Helmholtz time params $RL^{2}E$ MSEMSE Net
[…]